### v2

# library.bib

Abstract
We extend Friedman’s theorem to show that, for any fixed $r$, the random $2r$–regular Schreier graphs depicting the action of $r$ random permutations of $[n]$ on $k_{n}$–tuples of distinct elements in $[n]$ are a.a.s. weakly Ramanujan, for any $k_{n}\leq n^{\frac{1}{12}-\epsilon}.$ Previously this was known only for $k$–tuples where $k$ is fixed. In fact, we prove the stronger result of strong convergence of random matrices in irreducible representations of $S_{n}$ of quasi–exponential dimension.
We also give a new bound for the expected stable irreducible character of a random permutation obtained via a word map, showing that $\mathbb{E}\left[\chi^{\mu}\left(w(\sigma_{1},\dots,\sigma_{r})\right)\right]=O\left(n^{-k}\right)$, where $k$ is the number of boxes outside the first row of the Young diagram $\mu,$ solving one aspect of a conjecture of Hanany and Puder. We obtain this bound using an extension of Wise’s $w$–cycle conjecture.
1 Introduction
Let $G_{n}$ be a connected $2r$–regular graph on $n$ vertices, $A_{G_{n}}$ its adjacency matrix with eigenvalues $2r=\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{n}$. Let $\lambda(G_{n})=\mathrm{max}\left(\lambda_{2},|\lambda_{n}|\right).$ The Alon–Boppana bound dictates that
$\lambda\left(G_{n}\right)\geq 2\sqrt{2r-1}-o(1),$
so that $2r$–regular graphs with $\lambda\left(G_{n}\right)\leq 2\sqrt{2r-1}$ have optimal spectral gap, see [Alon86, Friedman03, Nilli91] for example.
Alon conjectured in [Alon86] that most $d$–regular graphs on a large number of vertices are weakly Ramanujan with high probability, and a breakthrough paper by Friedman [Friedman2008] proved Alon’s conjecture, proving that random $2r$–regular Schreier graphs (see §2.8), $\mathrm{Sch}\left(S_{n}\curvearrowright[n],\sigma_{1},\dots,\sigma_{r}\right)$ are asymptotically almost surely (a.a.s.) weakly Ramanujan, where a.a.s. means with probability $\to 1$ as $n\to\infty.$
Questions remain about the expansion properties of Cayley graphs of $S_{n}.$ These questions are very different in the sense that the Cayley graphs have $n!\sim n^{n}$ vertices, rather than only $n$ vertices. Kassabov, in [Kassabov], proved the following.
Theorem 1.1 (Kassabov).
There exist $L,\epsilon>0$ such that, for any $n$, there exists an explicit generating set $X_{n}$ of $S_{n}$, of size at most $L$, such that the Cayley graphs $\bigg{\{}\mathrm{Cay}\left(S_{n},X_{n}\right)\bigg{\}}_{n\geq 1}$ form a family of $\epsilon$–expanders.
Whether or not random Cayley graphs of $S_{n}$ (ie with a random generating set) are uniform expanders, or if they satisfy the conclusion of Friedman’s theorem, are still open questions. Our main result below is in this direction.
For any integers $r,k>0,$ define $\mathcal{G}_{r}\left(n,k\right)$ to be the collection of Schreier graphs $\mathrm{Sch}\left(S_{n}\curvearrowright[n]_{k},\sigma_{1},\dots,\sigma_{r}\right),$ where $[n]_{k}$ denotes the set of all $k$–tuples of distinct elements in $[n]$.
Theorem 1.2.
Fix any integer $r>1$ and let $\alpha<\frac{1}{12}.$ For any sequence of positive integers $\left(k_{n}\right)_{n\geq 1}$ with $k_{n}\leq n^{\alpha}$, let $\left(G_{n}\right)_{n\geq 1}$ be a sequence of random $2r$–regular Schreier graphs, where for each $n\geq 1$, $G_{n}\in\mathcal{G}_{r}\left(n,k_{n}\right)$ is obtained by choosing $r$ i.i.d. uniformly random permutations, $\sigma_{1},\dots,\sigma_{r}\in S_{n}.$
Then, for any $\epsilon>0,$
$\mathbb{P}\left[\lambda\left(G_{n}\right)\leq 2\sqrt{2r-1}+\epsilon\right]\overset{n\rightarrow\infty}{\longrightarrow}1.$
The case $\alpha=0$ is exactly Friedman’s theorem. The collection of Schreier graphs $\mathcal{G}_{r}\left(n,n\right)$ is the collection of Cayley graphs of $S_{n},$ so proving the above result for $\alpha=1$ would show that random Cayley graphs of $S_{n}$ are a.a.s. weakly Ramanujan. Nevertheless, Theorem 1.2 is the first result showing that the Schreier graphs constructed using random permutations acting on tuples of distinct elements have near–optimal spectral gap with high probability, in the case where the size of the tuple is allowed to grow with $n.$
Previously, this was known for tuples of size $1$ by [Friedman2008] (and later, with a shorter proof, by Bordenave [Bordenave2020]); for pairs of distinct elements by Bordenave and Collins [BordenaveCollins2018] and more recently, for $k$–tuples for any fixed $k$, by Chen, Garza–Vargas, Tropp and van Handel in [CGVTvH2024].
Remark 1.3.
Prior to the result for fixed $k$ in [CGVTvH2024], it was already known that for any fixed $k,$ a random regular Schreier graph of $S_{n}$ has a uniform spectral gap with high probability. Indeed, in [Friedmanetal, Theorem 2.1], they prove that, for any fixed $k,r>0,$ if $G_{n}\in\mathcal{G}_{r}\left(n,k\right)$ is a random Schreier graph, then for any $\epsilon>0,$
$\mathbb{P}\left[\lambda\left(G_{n}\right)\leq(1+\epsilon)2r\left(\frac{\sqrt{2r-1}}{r}\right)^{\frac{1}{r+1}}\right]\overset{n\to\infty}{\longrightarrow}1.$
Remark 1.4.
A comment on ‘derandomization’. In [Mohanty], using Bordenave’s proof of Friedman’s theorem, they give a construction of a regular graph on $O(N)$ vertices, that is near–Ramanujan with high probability (probability $\geq 1-O\left(\frac{\log N}{2^{100\left(\log N\right)^{\frac{1}{4}}}}\right)$). They obtain such a graph by randomly choosing a ‘seed’ $s\in\{0,1\}^{O\left(\log(N)^{2}\right)}$ and using this as the input to an algorithm. That is, their construction requires $\sim\log(N)^{2}$ random bits. With the construction in Theorem 1.2, one can construct a regular graph on $N$ vertices that is near–Ramanujan with high probability (probability $\geq 1-O\left(\frac{1}{\left(\log\left(N\right)^{12}\right)^{1-(12+\epsilon)\alpha}}\right)$), using $\sim\log(N)^{12}$ random bits, by sampling random permutations $\sigma_{1},\dots,\sigma_{r}\in S_{n}$, with $n\sim\log(N)^{12}$. Although this requires slightly more random bits than the algorithm given in [Mohanty], it is a more algebraic construction and illustrates a different approach to such a problem.  [^1] We thank Ramon van Handel for pointing out this application.
Our proof of Theorem 1.2 relies on the remarkable new approach to strong convergence detailed in [CGVTvH2024], see §1.1, as well as the additional criterion for temperedness of arbitrary functions on finitely generated groups $\Gamma$ with a finite fixed generating set (an adaptation of the classical notion of a tempered representation) given by Magee and de la Salle in [MAgeeDelaSalle].
To use this approach effectively, we prove our other main theorem pertaining to word maps and stable irreducible characters (see §2.9), in which we prove a new bound on the expected stable irreducible character of a $w$–random permutation. Given a Young diagram $\lambda\vdash k$ and $n\geq k+\lambda_{1},$ we define the Young diagram $\lambda^{+}(n)=\left(n-k,\lambda\right)\vdash n$, see §2.1.
Theorem 1.5.
Let $w\in F_{r}=\langle x_{1},\dots,x_{r}\rangle$ be a word in the free group on $r$ generators with $w\neq e$. Let $k\in\mathbb{Z}_{>0}$ and $\lambda\vdash k$. Then $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ is a rational expression in $n$ and, if $w$ is not a proper power, then
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=O\left(\frac{1}{n^{k}}\right).$
Remark 1.6.
That $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ can be written as a rational expression in $n$ follows from (24) or indeed by combining [Nica, LinialPuder] with [HananyPuder, Proposition B.2]. In the case of $w$ being a proper power, the conclusion of Theorem 1.5 follows from [Nica] and [LinialPuder, Section 4].
Remark 1.7.
We prove Theorem 1.5 using ‘combinatorial integration’. We use a projection formula obtained in our previous paper [Cassidy2023] to translate the computation of $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ in to combinatorial problem. We then prove an extension of Wise’s $w$–cycles conjecture (see [LouderWilton, HelferWise, Wise]), a theorem in low–dimensional topology, to obtain our final bound.
Theorem 1.5 generalizes the work of Nica [Nica] and Linial and Puder [LinialPuder], where it is established that, if #$\#\mathrm{fix}(\sigma)$ is the number of fixed points of a permutation $\sigma\in S_{n},$ then, if $w$ is not a proper power,
$\mathbb{E}_{w}\left[\#\mathrm{fix}(\sigma)-1\right]=O\left(\frac{1}{n}\right)$ (1)
Of course, #$\#\mathrm{fix}-1$ is the character of the standard representation of $S_{n},$ $\chi^{(n-1,1)}$, so that this is a special case of our theorem. In [PuderParzanchewski2012], Puder and Parzanchevski improve this, showing that
$\mathbb{E}_{w}\left[\#\mathrm{fix}(\sigma)-1\right]=O\left(\frac{1}{n^{\pi(w)-1}}\right)$
where $\pi(w)$ is the primitivity rank of the word $w$ (see, for example [PuderParzanchewski2012, Definition 1.7], for the definition). Hanany and Puder [HananyPuder, Theorem 1.3] generalize this result to other stable irreducible representations of $S_{n}.$ They show that, if $w$ is not a proper power, then for any $k\neq 0,1$ and for any $\lambda\vdash k,$
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=O\left(\frac{1}{n^{\pi(w)}}\right).$
Theorem 1.5 is thus an improvement on this bound in the regime $k>\pi(w).$
In [HananyPuder, Conjecture 1.8], it is conjectured that
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=O\left(\frac{1}{n^{k(\pi(w)-1)}}\right).$
So Theorem 1.5 confirms the conjecture for words $w$ with primitivity rank $\pi(w)=2.$ In cases where there is no further information on $w$ other than $w\neq e$ and $w$ is not a proper power, the bound given in Theorem 1.5 is expected to be sharp. In fact, it has recently been conjectured by Puder and Shomroni (see [PuderShomroni, Conjecture 1.2]) that we should have
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=O\left(\frac{1}{n^{k(\mathrm{s}\pi(w))}}\right),$
where $\mathrm{s}\pi(w)$ is a property of $w$ known as the stable primitvity rank (this was first introduced by Wilton in [Wilton], indeed see Definition 10.6 in (ibid.) or [PuderShomroni, Definition 4.1] for an equivalent definition). They also conjecture that this bound is the best one possible. That is, for every word $w\in F_{r}$, for every $k,$ there is a Young diagram $\lambda\vdash k$ such that $\frac{1}{n^{k(\mathrm{s}\pi(w))}}=O\left(\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]\right).$ An additional conjecture of Wilton (see eg [PuderShomroni, Conjecture 4.7]) posits that, for any $w\in F_{r},$ we have the equality $\mathrm{s}\pi(w)=\pi(w)-1$, so that the conjecture of Hanany and Puder is a combination of these two conjectures.
1.1 Strong Convergence
As we noted earlier, we actually prove a stronger statement than Theorem 1.2, that of strong convergence of random representations of $F_{r}=\left\langle x_{1},\dots,x_{r}\right\rangle$ that factor through permutations.
Definition 1.8.
Given a sequence of random, finite dimensional, unitary representations $\{\pi_{n}:F_{r}\to U\left(N_{n}\right)\}_{n\geq 1},$ we say that $\pi_{n}$ strongly converge to the left regular representation $\lambda:F_{r}\to U\left(l^{2}\left(F_{r}\right)\right)$ a.a.s. if, for any $z\in\mathbb{C}\left[F_{r}\right]$, for any $\epsilon>0$, we have
$\mathbb{P}\left[\left|\left\|\pi_{n}(z)\right\|-\left\|\lambda(z)\right\|\right|>\epsilon\right]\overset{n\to\infty}{\longrightarrow}1.$ (2)
The norms here are operator norms, see §2.6.
We are interested in the cases whereby the representations $\pi_{n}=\rho\circ\phi_{n},$ where $\phi_{n}\in\mathrm{Hom}\left(F_{r},S_{n}\right)$ is random and $\rho:S_{n}\to\mathrm{End}\left(V\right)$ is a representation of $S_{n}$ of dimension $N_{n}.$ The power of strong convergence in this setting is that, since the convergence must hold for every $z\in\mathbb{C}\left[F_{r}\right]$, it allows us to prove results like Theorem 1.2, which require (2) to hold only for specific elements of $\mathbb{C}\left[F_{r}\right].$
For example, proving Friedman’s Theorem is equivalent to showing that
$\mathbb{P}\left[\left|\left\|\pi_{n}(z)\right\|-2\sqrt{2r-1}\right|>\epsilon\right]\overset{n\to\infty}{\longrightarrow}1$ (3)
where
$z=x_{1}+x_{1}^{-1}+\dots+x_{r}+x_{r}^{-1}$ (4)
and $\pi_{n}=\rho^{(n-1,1)}\circ\phi_{n}$, with $\phi_{n}\in\mathrm{Hom}\left(F_{r},S_{n}\right)$ uniformly random (ie obtained by uniformly randomly choosing a permutation $\sigma_{i}\in S_{n}$ for each $i$ and mapping $x_{i}\mapsto\sigma_{i}$) and where
$\rho^{(n-1,1)}:\mathbb{C}\left[S_{n}\right]\to\mathrm{End}\left(V^{(n-1,1)}\right)$
is the standard representation of $S_{n}.$ This is because, if $\rho:S_{n}\to\mathrm{End}\left(\mathbb{C}^{n}\right)$ is the permutation representation of $S_{n},$ then
$\rho\left(\sigma_{1}+\sigma_{1}^{-1}+\dots+\sigma_{r}+\sigma_{r}^{-1}\right)$
is the adjacency matrix of the graph $\mathrm{Sch}\left(S_{n}\curvearrowright[n],\sigma_{1},\dots,\sigma_{r}\right)$ and $2\sqrt{2r-1}=\|\lambda(z)\|$ in this case. We consider the orthogonal complement to the $S_{n}$–invariant vectors in $\mathbb{C}^{n}$ to account for the trivial eigenvalue $2r.$ In the case of $\rho$ above, that is the standard representation.
More generally, for any $k\in\mathbb{Z}_{>0}$ we define
$\bar{\rho}_{n,k}:\mathbb{C}\left[S_{n}\right]\to\mathrm{End}\left(\left(\mathbb{C}^{n}\right)^{\otimes k}\right),$
the $k^{\mathrm{th}}$ tensor power of the permutation representation. We then define
$\rho_{n,k}:\mathbb{C}\left[S_{n}\right]\to\mathrm{End}\left(V_{n,k}\right)$
to be the restriction of $\bar{\rho}_{n,k}$ to the orthocomplement to the $S_{n}$–invariant vectors, $V_{n,k}\subseteq\left(\mathbb{C}^{n}\right)^{\otimes k}$ and define a random sequence of unitary representations of $F_{r},$$\{\pi_{n,k}\overset{\mathrm{def}}{=}\rho_{n,k}\circ\phi_{n}\}_{n\geq 1},$ with $\phi_{n}\in\mathrm{Hom}\left(F_{r},S_{n}\right)$ uniformly random as before.
Bordenave and Collins [BordenaveCollins2018] proved strong convergence to the left regular representation $\lambda:F_{r}\to U\left(l^{2}\left(F_{r}\right)\right)$ a.a.s. for the sequence of random representations of $F_{r}$, $\{\pi_{n,1}\}_{n\geq 1}$. They extend their results to show strong convergence a.a.s. for $\{\pi_{n,2}\}_{n\geq 1}$. With the new approach in [CGVTvH2024], strong convergence a.a.s. for $\{\pi_{n,k}\}_{n\geq 1}$ for any fixed $k$ is now known. We adopt and extend this new approach to prove our final main theorem below.
Theorem 1.9.
For any $\alpha<\frac{1}{12}$, for any $z\in\mathbb{C}\left[F_{r}\right]$ and for any $\epsilon>0,$
$\mathbb{P}\left[\sup_{k_{n}\leq n^{\alpha}}\left|\left\|\pi_{n,k_{n}}(z)\right\|-\left\|\lambda(z)\right\|\right|<\epsilon\right]\overset{n\to\infty}{\longrightarrow}1.$
Remark 1.10.
Theorem 1.2 follows immediately from Theorem 1.9 by taking $z=x_{1}+x_{1}^{-1}+\dots+x_{r}+x_{r}^{-1}.$ Theorem 1.9 is a much stronger statement, and it is certainly not necessary prove such a statement to prove Theorem 1.2.
1.2 Overview of paper
§2 gives background information on variations of Schur–Weyl duality, including a refinement of Schur–Weyl–Jones duality and a projection formula given by the author in [Cassidy2023]. We also give a brief overview of the Weingarten calculus for the symmetric group and a number of necessary results from [MAgeeDelaSalle].
§3 contains our proof of Theorem 1.9, assuming Theorem 1.5 and Theorem 4.1. The proofs of Theorem 1.5 and Theorem 4.1 are given in §4.
Acknowledgements
This paper is part of a project that received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 949143).
We thank Michael Magee for the many conversations regarding this paper, as well as Doron Puder, Lars Louder and Ramon van Handel for their valuable input. We also thank Noam Ta Shma for his comments and feedback on our initial version of this paper.
2 Background
A note on notation
We will use the following Vinogradov notation. Given two functions $f,g$ of $n\in\mathbb{Z}_{>0},$ we will write $f\ll g$ to indicate that that there exist constants $C$ and $N$ such that $f(n)\leq Cg(n)$ for all $n\geq N.$ We will then write $f=O(g)$ to mean that $f\ll|g|$ and $f\asymp g$ to mean that $f\ll g$ and $f\ll g.$ If the implied constants depend on some additional parameter, then these will be added as subscripts.
2.1 Symmetric group
We will denote by $S_{n}$ the symmetric group on $n$ elements and view this as the group of permutations of $[n]\overset{\mathrm{def}}{=}\{1,\dots,n\}$.
A Young diagram (YD) $\mu=\left(\mu_{1},\dots,\mu_{l(\mu)}\right)\vdash n$ is an arrangement of boxes in to aligned rows, in which row $i$ contains $\mu_{i}$ boxes. By definition, we have $\mu_{1}\geq\mu_{2}\geq\dots\geq\mu_{l(\mu)}>0,$ so that the number of boxes in each row is non–increasing as the row index increases. We also have $\mu_{1}+\dots+\mu_{l(\mu)}=|\mu|=n.$
The equivalence classes of complex irreducible representations of $S_{n}$ are in a natural bijection with the set of Young diagrams $\mu\vdash n,$ see [VershikOkounkov] for example. We will denote the representation corresponding to $\mu$ by $V^{\mu}$ and denote its character by $\chi^{\mu}.$
Every finite–dimensional complex representation $V$ of $S_{n}$ admits a decomposition
$V\cong\bigoplus_{\mu\vdash n}\left(V^{\mu}\right)^{\oplus a_{\mu}},$
unique up to isomorphism (ie unique up to isomorphisms that preserve each component $\left(V^{\mu}\right)^{\oplus a_{\mu}}$ of the decomposition). For each $\mu\vdash n,$ the subspace $\left(V^{\mu}\right)^{\oplus a_{\mu}}$ is called the $V^{\mu}$–isotypic component of $V.$
We will denote by $d_{\mu}\overset{\mathrm{def}}{=}\chi^{\mu}(e),$ so that $d_{\mu}=\mathrm{dim}\left(V^{\mu}\right).$ Given a representation $V$ of $S_{n}$ and a permutation $\sigma,$ we may write $V(\sigma)$ or just $\sigma$ to mean the element of $\mathrm{End}\left(V\right)$ corresponding to $\sigma,$ defined by the representation.
Given $\lambda=(\lambda_{1},\dots,\lambda_{l(\lambda)}),$ , we define $V^{\lambda^{+}(n)}$ to be the irreducible representation of $S_{n}$ (when $n-|\lambda|>\lambda_{1}$) corresponding to the YD
$\lambda^{+}(n)=\left(n-|\lambda|,\lambda_{1},\dots,\lambda_{l(\lambda)}\right)\vdash n.$
2.2 Partition algebra
The partition algebra appeared in the work of Martin [Martin1994], Martin and Saleur [MartinSaleur1992] and Jones [Jones] as a generalization of the Temperley–Lieb algebra and the Potts model in statistical mechanics. It is closely connected to the symmetric group via Schur–Weyl–Jones duality, which was first detailed explicitly by Jones in [Jones].
2.2.1 Set partitions
A set partition of the set $[m]$ is a grouping of the elements in to non–empty, disjoint subsets in which every element is in exactly one subset. We will write $i\sim j$ to indicate that $i$ and $j$ are in the same subset of a partition, and refer to the subsets as blocks. The number of blocks of a partition $\pi\in\mathrm{Part}\left([m]\right)$ will be denoted $|\pi|.$ Obviously, $1\leq|\pi|\leq m$.
We denote by $\mathrm{Part}\left([m]\right)$ the set of set partitions of $[m]$. One may define a natural partial ordering on $\mathrm{Part}\left([m]\right)$ by refinement: $\pi_{1}\leq\pi_{2}$ if every block of $\pi_{1}$ is contained in a block of $\pi_{2}.$ We say that $\pi_{1}$ is finer than $\pi_{2}$ or that $\pi_{2}$ is coarser than $\pi_{1}.$
From this partial ordering one may derive the Möbius function $\mu$ on $\mathrm{Part}\left([m]\right)$. If $\pi_{2}=\Big{\{}\mathcal{P}_{1},\dots,\mathcal{P}_{|\pi_{2}|}\Big{\}}$ and $\pi_{1}$ is a refinement of $\pi_{2}$ in which each block $\mathcal{P}_{i}$ is split in to $b_{i}$ subsets, then
$\mu\left(\pi_{1},\pi_{2}\right)=(-1)^{|\pi_{1}|-|\pi_{2}|}\prod_{i=1}^{|\pi_{2}|}(b_{i}-1)!,$
as is described in the foundational paper by Rota [Rota].
Given partitions $\pi_{1},\pi_{2}\in\mathrm{Part}\left([m]\right),$ their meet $\pi_{1}\wedge\pi_{2}$ is defined by $i\sim j$ in $\pi_{1}\wedge\pi_{2}$$\iff i\sim j$ in both $\pi_{1}$ and $\pi_{2}.$ Their join $\pi_{1}\vee\pi_{2}$ is defined by $i\sim j$ in $\pi_{1}\vee\pi_{2}$$\iff i\sim j$ in either $\pi_{1}$ or $\pi_{2}.$
2.2.2 Partition algebra
Given $k\in\mathbb{Z}_{>0}$ and $n\in\mathbb{C}$, the partition algebra $P_{k}(n)$ is an associative unital algebra over $\mathbb{C}.$ It is the vector space
$\big{\langle}\pi:\ \pi\in\mathrm{Part}\left([2k]\right)\big{\rangle}_{\mathbb{C}},$
with a natural multiplication as described below
To each $\pi\in\mathrm{Part}\left([2k]\right)$ we associate a diagram consisting of two rows of $k$ vertices labeled from $1,\dots k$ on the top row and from $k+1,\dots,2k$ on the bottom row. We draw an edge between any two vertices $i$ and $j$ if and only if $i\sim j$ in $\pi.$
Any diagram consisting of two rows of $k$ vertices with all connected components complete clearly corresponds to an element $\pi\in\mathrm{Part}\left([2k]\right).$
The product $\pi_{1}\pi_{2}$ is obtained by merging the bottom row of vertices of $\pi_{1}$ with the top row of vertices of $\pi_{2}$ to obtain a new diagram $\tilde{\pi}_{3}$ with three rows of $k$ vertices; adding edges between any vertices that are in the same connected component (where there is not already an edge); removing the middle row of vertices and any adjacent edges to obtain $\pi_{3}\in\mathrm{Part}\left([2k]\right).$ Then we have
$\pi_{1}\pi_{2}=n^{\gamma}\pi_{3},$
where $\gamma$ is the number of connected components of $\tilde{\pi}_{3}$ with vertices solely in the middle row.
Martin and Saleur [MartinSaleur1992] describe some important properties of $P_{k}(n).$
Theorem 2.1.
The partition algebra $P_{k}(n)$ is semisimple for any $n\in\mathbb{C},$ unless $n$ is an integer in the range $[0,2k-1].$
Theorem 2.2.
If $P_{k}(n)$ is semisimple, then there is a natural bijection between the equivalence classes of simple modules over $P_{k}(n)$ and the set
$\Lambda_{k,n}\overset{\mathrm{def}}{=}\{\lambda\vdash i:\ i=0,\dots,k\}.$
By convention, there is exactly one Young diagram $\lambda\vdash 0.$ We will write $R^{\lambda}$ to denote the equivalence class corresponding to $\lambda.$
2.3 Schur–Weyl–Jones duality
There is an analogue of classical Schur–Weyl duality due to Jones [Jones], which we refer to as Schur–Weyl–Jones duality.
There is a natural right action of $P_{k}(n)$ on the vector space $\left(\mathbb{C}^{n}\right)^{\otimes k}.$ For any standard orthonormal basis vectors $e_{i_{i}}\otimes\dots\otimes e_{i_{k}}$ and $e_{i_{k+1}}\otimes\dots\otimes e_{i_{2k}}$ and any $\pi\in\mathrm{Part}\left([2k]\right),$ we have
$\big{\langle}\left(e_{i_{i}}\otimes\dots\otimes e_{i_{k}}\right)\pi,\ e_{i_{k+1}}\otimes\dots\otimes e_{i_{2k}}\big{\rangle}=\begin{cases}1&\mathrm{if}\ j\sim l\mathrm{\ in}\ \pi\implies i_{j}=i_{l}\\ 0&\mathrm{otherwise.}\end{cases}$
This extends linearly to define a right action on $\left(\mathbb{C}^{n}\right)^{\otimes k}.$ For each $\pi\in P_{k}(n),$ we define $p_{\pi}\in\mathrm{End}\left(\left(\mathbb{C}^{n}\right)^{\otimes k}\right),$ where $p_{\pi}(v)=v\pi.$
The symmetric group $S_{n}$ acts on $\left(\mathbb{C}^{n}\right)^{\otimes k}$ on the left via the diagonal permutation action, ie for $g\in S_{n}$ and basis vector $e_{i_{i}}\otimes\dots\otimes e_{i_{k}}\in\left(\mathbb{C}^{n}\right)^{\otimes k},$
$g\left(e_{i_{i}}\otimes\dots\otimes e_{i_{k}}\right)=e_{g\left(i_{1}\right)}\otimes\dots\otimes e_{g\left(i_{k}\right)}.$
Schur–Weyl–Jones duality asserts the following theorem.
Theorem 2.3.
When $S_{n}$ and $P_{k}(n)$ act as above and $n\geq 2k$,
  1. 1.
    $P_{k}(n)$ generates $\mathrm{End}_{S_{n}}\left(\left(\mathbb{C}^{n}\right)^{\otimes k}\right)$,
  2. 2.
    $\mathbb{C}\left[S_{n}\right]$ generates $\mathrm{End}_{P_{k}(n)}\left(\left(\mathbb{C}^{n}\right)^{\otimes k}\right)$.
  3. 3.
    As a $\left(S_{n},P_{k}(n)\right)$–bimodule,
    $\left(\mathbb{C}^{n}\right)^{\otimes k}\cong\bigoplus_{\lambda\in\Lambda_{k,n}}V^{\lambda^{+}(n)}\otimes R^{\lambda}.$
2.4 Refinement
In [Cassidy2023], the author gives a refinement of Schur–Weyl–Jones duality. We give a brief overview of the main results of this paper. We will assume that $k\in\mathbb{Z}_{>0}$ and $n\geq 2k$.
There is an algebra inclusion
$\iota:\mathbb{C}\left[S_{k}\right]\to P_{k}(n),$
where, for each $\sigma\in S_{k},$
$\iota\left(\sigma\right)=\big{\{}\{1,k+\sigma^{-1}(1)\},\dots,\{k,k+\sigma^{-1}(k)\}\big{\}}.$
There is also a corresponding restriction map
$R:P_{k}(n)\to\mathbb{C}\left[S_{k}\right]$
in the sense that $R\circ\iota=\mathrm{Id}:\mathbb{C}\left[S_{k}\right]\to\mathbb{C}\left[S_{k}\right]$ with $R(\pi)=0$ for any set partition $\pi\in\mathrm{Part}\left([2k]\right)$ that is not the image of some permutation in $S_{k}.$
Notice that the image of $\iota$ is the complex linear span of all partitions that contain exactly one element $i$ with $0\leq i\leq k$ and exactly one element $j$ satisfying $k+1\leq j\leq 2k$ (equivalently, all partitions for which the corresponding diagram consists of $k$ connected components, each containing exactly one vertex from each row).
Definition 2.4.
We define the linear contraction map $T_{j}:\left(\mathbb{C}^{n}\right)^{\otimes k}\to\left(\mathbb{C}^{n}\right)^{\otimes k-1}$ by
$T_{j}\left(e_{i_{i}}\otimes\dots\otimes e_{i_{k}}\right)$
$=$ $e_{i_{1}}\otimes\dots\otimes\dot{e}_{i_{j}}\otimes\dots\otimes e_{i_{k}}$
=$\overset{\mathrm{def}}{=}$ $e_{i_{1}}\otimes\dots\otimes e_{i_{j-1}}\otimes e_{i_{j+1}}\otimes\dots\otimes e_{i_{k}}.$
Definition 2.5.
The vector subspace $D_{k}(n)\subset\left(\mathbb{C}^{n}\right)^{\otimes k}$ is defined to be
$D_{k}(n)\overset{\mathrm{def}}{=}\left\langle e_{i_{1}}\otimes\dots\otimes e_{i_{k}}:|\{i_{1},\dots,i_{k}\}|=k\right\rangle_{\mathbb{C}},$
the complex linear span of all $k$–tuples of distinct elements.
We use the contraction maps and $D_{k}(n)$ to construct a subspace on which we have a version of Schur–Weyl duality involving only the symmetric groups.
Definition 2.6.
The vector subspace $A_{k}(n)\subseteq\left(\mathbb{C}^{n}\right)^{\otimes k}$ is defined by
$A_{k}(n)=D_{k}(n)\cap\bigcap_{j=1}^{k}\mathrm{Ker}\left(T_{j}\right).$
Lemma 2.7.
We have
$A_{k}(n)=\bigcap_{\pi\in\mathrm{Ker}\left(R\right)}\mathrm{Ker}\left(\pi\right).$
Which is to say that, for $\pi\in\mathrm{Part}\left([2k]\right)$ and $v\in A_{k}(n),$ we have $p_{\pi}(v)=0$ unless $\pi=\iota(\sigma)$ for some $\sigma\in S_{k}.$ Moreover, if $\sigma\in S_{k},$ then $p_{\iota(\sigma)}\in\mathrm{End}\left(\left(\mathbb{C}^{n}\right)^{\otimes k}\right)$ permutes the tensor coordinates according to $\sigma.$ That is,
$p_{\iota(\sigma)}\left(e_{i_{i}}\otimes\dots\otimes e_{i_{k}}\right)=e_{i_{\sigma(1)}}\otimes\dots\otimes e_{i_{\sigma(k)}}.$
This extends linearly to define an action of $\mathbb{C}\left[S_{k}\right]$ on $\left(\mathbb{C}^{n}\right)^{\otimes k}$ (and suitable subspaces).
We then view $A_{k}(n)$ as a representation of $S_{n}\times S_{k},$ where once again $g\in S_{n}$ acts via the diagonal permutation action and $\sigma\in S_{k}$ acts via $p_{\iota\left(\sigma^{-1}\right)}.$ We denote this representation
$\Delta:\mathbb{C}\left[S_{n}\times S_{k}\right]\to\mathrm{End}\left(A_{k}(n)\right).$
The version of Schur–Weyl duality on $A_{k}(n)$ is given in the following theorem from [Cassidy2023].
Theorem 2.8.
Where $S_{n}$ and $S_{k}$ act as above,
  1. 1.
    $\mathbb{C}\left[S_{k}\right]$ generates $\mathrm{End}_{S_{n}}\left(A_{k}(n)\right)$,
  2. 2.
    $\mathbb{C}\left[S_{n}\right]$ generates $\mathrm{End}_{S_{k}}\left(A_{k}(n)\right)$.
  3. 3.
    The decomposition of $A_{k}(n)$ in to irreducible $S_{n}\times S_{k}$ representations is
    $A_{k}(n)\cong\bigoplus_{\lambda\vdash k}V^{\lambda^{+}(n)}\otimes V^{\lambda}.$
The main result of [Cassidy2023] is a formula for the orthogonal projection from $\left(\mathbb{C}^{n}\right)^{\otimes k}$ to subspaces isomorphic to the irreducible ‘blocks’ in the decomposition of $A_{k}(n).$ Define
$\xi\overset{\mathrm{def}}{=}\left(e_{1}-e_{2}\right)\otimes\dots\otimes\left(e_{2k-1}-e_{2k}\right)\in A_{k}(n).$
We denote by $\xi_{\lambda}^{\mathrm{norm}}$ the normalised projection of $\xi$ to the $V^{\lambda}$–isotypic subspace of $A_{k}(n)$. This can be shown to be non–zero, giving the following lemma.
Lemma 2.9.
For any $\lambda\vdash k$, the $S_{n}\times S_{k}$ representation generated by $\xi_{\lambda}^{\mathrm{norm}}$ is isomorphic to $V^{\lambda^{+}(n)}\otimes V^{\lambda}.$ That is,
$\mathcal{U}_{\lambda,n}\overset{\mathrm{def}}{=}\big{\langle}\Delta\left(g,\sigma\right)\left(\xi_{\lambda}^{\mathrm{norm}}\right):\ g\in S_{n},\ \sigma\in S_{k}\big{\rangle}_{\mathbb{C}}\cong V^{\lambda^{+}(n)}\otimes V^{\lambda}.$
This leads to the main theorem and a corollary of the author in [Cassidy2023] which will be key for the combinatorial integration method used in this paper.
Theorem 2.10.
For any $\lambda\vdash k$, the orthogonal projection
$\mathcal{Q}_{\lambda,n}:\left(\mathbb{C}^{n}\right)^{\otimes k}\to\mathcal{U}_{\lambda,n}$
is given by
$\sum_{\pi\in\mathrm{Part}\left([2k]\right)}c(n,k,\lambda,\pi)p_{\pi},$
where
$c(n,k,\lambda,\pi)=d_{\lambda}d_{\lambda^{+}(n)}\frac{(-1)^{k}}{k!}\left(\sum_{\sigma,\sigma^{\prime}\in S_{k}}\sum_{\pi^{\prime}\leq\iota\left(\sigma^{-1}\sigma^{\prime}\right)\wedge\pi}\chi^{\lambda}(\sigma)\chi^{\lambda}(\sigma^{\prime})\frac{(-1)^{|\pi^{\prime}|}\mu\left(\pi^{\prime},\pi\right)}{(n)_{|\pi^{\prime}|}}\right).$ (5)
Corollary 2.11.
For any $\lambda\vdash k$ and any $g\in S_{n},$ $\chi^{\lambda^{+}(n)}(g)$ is equal to
$\frac{1}{d_{\lambda}}\mathrm{tr}_{\mathcal{U}_{\lambda,n}}\left(g\right)$
$=$ $\frac{1}{d_{\lambda}}\mathrm{btr}_{\left(\mathbb{C}^{n}\right)^{\otimes k}}\left(g,\mathcal{Q}_{\lambda,n}\right).$
If $\pi\leq\iota(\tau)$ for some $\tau\in S_{k},$ then the diagram corresponding to $\pi$ is obtained from that of $\iota(\tau)$ by deleting edges. Let $\mathrm{del}\left(\pi\right)$ be the number of edges deleted (ie $\mathrm{del}\left(\pi\right)=|\pi|-k$, so that if $\pi\leq\iota\left(\tau_{1}\right)$ and $\pi\leq\iota\left(\tau_{2}\right),$ then $\mathrm{del}(\pi)$ is independent of whether we count the edges deleted from $\tau_{1}$ or $\tau_{2}$).
Lemma 2.12.
If $\pi\leq\iota(\tau)$ for some $\tau\in S_{k},$ then
$c(n,k,\lambda,\pi)=O\left(\frac{1}{n^{\mathrm{del}(\pi)}}\right).$
Proof.
The terms in (5) with the greatest power of $n$ occur when $\pi^{\prime}=\pi.$ In this case, we have some coefficient multiplied by
$n^{k}n^{-|\pi|}=n^{k-|\pi|}=n^{-\mathrm{del}(\pi)}.$
Remark 2.13.
If $\pi=\iota(\tau)$, then this is $O(1).$
2.5 The Weingarten calculus
An important part of the proof of Theorem 2.10 is the Weingarten calculus. Originally developed by Weingarten [Weingarten] and Xu [Xu] and then more rigorously by Collins [Collins2003] and Collins and Sniady [CollinsSniady2006], the Weingarten calculus is a method for computing integrals of matrix coefficients using Schur–Weyl duality analogues.
Below is the main formulation for the Weingarten calculus for the symmetric group, we direct the reader to the survey [CollinsWeingartenShort, Section 3] or indeed any of the aforementioned papers for an in depth discussion.
Define, for any multi–index $I=\left(i_{1},\dots,i_{m}\right)$ with $i_{1},\dots,i_{m}\in[n]$ and any set partition $\pi\in\mathrm{Part}\left([m]\right)$ , the function $\delta_{\pi},$ where
$\delta_{\pi}\left(I\right)=\begin{cases}1&\mathrm{if}\ j\sim l\mathrm{\ in}\ \pi\implies i_{j}=i_{l}\\ 0&\mathrm{otherwise}.\end{cases}$
Theorem 2.14.
For any multi–indices $I=\left(i_{1},\dots,i_{m}\right),$ $J=\left(j_{1},\dots,j_{m}\right)\in[n]^{m},$ with respect to the uniform measure, we have
$\mathop{\int}_{S_{n}}g_{i_{1}j_{1}}\dots g_{i_{m}j_{m}}dg=\sum_{\pi_{1},\pi_{2}\in\mathrm{Part}\left([m]\right)}\delta_{\pi_{1}}(I)\delta_{\pi_{2}}(J)\mathrm{Wg}_{n,m}\left(\pi_{1},\pi_{2}\right),$
where $\mathrm{Wg}_{n,m}$ is the Weingarten function for $S_{n},$
$\mathrm{Wg}_{n,m}\left(\pi_{1},\pi_{2}\right)=\sum_{\pi\leq\pi_{1}\wedge\pi_{2}}\mu\left(\pi,\pi_{1}\right)\mu\left(\pi,\pi_{2}\right)\frac{1}{(n)_{|\pi|}}.$
Remark 2.15.
For any $\pi_{1},\pi_{2}\in\mathrm{Part}\left([m]\right),$ $\mathrm{Wg}_{n,m}\left(\pi_{1},\pi_{2}\right)=O\left(\frac{1}{n^{|\pi_{1}\wedge\pi_{2}|}}\right)$. In particular, $\mathrm{Wg}_{n,m}\left(\pi_{1},\pi_{1}\right)=O\left(\frac{1}{n^{|\pi_{1}|}}\right)$.
2.6 $C^{*}$–algebras of free groups
We refer to [NicaSpeicher] for a thorough introduction. To the free group $F_{r}$ we associate its group algebra $\mathbb{C}\left[F_{r}\right],$ the collection of elements of the form
$x=\sum_{w\in F_{r}}x(w)w,$
where each $x(w)\in\mathbb{C}$ and only finitely many are non–zero.
We define a norm on $\mathbb{C}\left[F_{r}\right]$ by
$\|x\|_{C^{*}\left(F_{r}\right)}=\sup\big{\{}\|\pi(x)\|_{\mathrm{op}}:\ \pi\ \mathrm{a\ representation\ of}\ F_{r}\big{\}},$
where $\|-\|_{\mathrm{op}}$ is the operator norm on $l^{2}\left(F_{r}\right)$. The completion of $\mathbb{C}\left[F_{r}\right]$ in this norm is the $C^{*}$–algebra of $F_{r},$ denoted $C^{*}\left(F_{r}\right)$.
The left regular representation of $F_{r}$ is denoted $\lambda$ and admits the reduced norm on $\mathbb{C}\left[F_{r}\right]$ by
$\|x\|_{\mathrm{C_{\lambda}^{*}\left(F_{r}\right)}}=\|\lambda\left(x\right)\|_{\mathrm{op}}.$
The completion of $\mathbb{C}\left[F_{r}\right]$ in this norm is the reduced $C^{*}$–algebra of $F_{r},$ denoted $C_{\lambda}^{*}\left(F_{r}\right).$
We define a trace $\tau$ on $\mathbb{C}\left[F_{r}\right]$ to be the map
$\tau(x)=x(e).$
This extends continuously to both the $C^{*}$–algebras defined above and remains a trace in both cases.
2.7 Polynomials and random walks
We collect some results and definitions on polynomials and random walks that are required for the analysis in §3. The proofs are readily available in [MAgeeDelaSalle]. See also [CGVTvH2024] for more background.
Lemma 2.16 ([MAgeeDelaSalle, Lemma 4.2]).
For every polynomial $P$ in one variable with bounded degree, $\deg\left(P\right)\leq D$ and for every integer $k\leq D,$
$\sup_{\left[0,\frac{1}{2D^{2}}\right]}\left|P^{(k)}\right|\leq\frac{2^{2k+1}D^{4k}}{(2k-1)!!}\sup_{n\geq D^{2}}\left|P\left(\frac{1}{n}\right)\right|,$
where $(2k-1)!!=(2k-1)(2k-1)\dots(3)1.$
2.7.1 Random walks and free groups
Definition 2.17.
We call a symmetric (ie $\mu(g)=\mu\left(g^{-1}\right))$ probability measure $\mu$ on the free group $F_{r}$ reasonable if its support is finite, contains the identity element and generates $F_{r}.$
We denote by $(g_{n})_{n\geq 0}$ the corresponding random walk in $F_{r}.$ So we write
$g_{n}=s_{1}\dots s_{n}$
where $s_{i}\in F_{r}$ are i.i.d. according to $\mu.$
The spectral radius $\rho=\rho(\mu)$ measures how fast the probability that a random walk returns to where it started decays. It is equal to the norm of $\lambda\left(\mu\right)$ on $l^{2}\left(F_{r}\right)$, where $\lambda$ is the left regular representation of $F_{r}.$
Define a proper power in the free group to be an element of the form $u^{d}$ for $u\in F_{r}$ and $d\geq 2.$
Proposition 2.18 ([MAgeeDelaSalle, Proposition 6.1]).
For any reasonable probability measure $\mu,$ there is a constant $C=C_{\mu}$ such that
$\mathbb{P}\left(g_{n}\ \mathrm{is\ a\ proper\ power}\right)\leq Cn^{5}\rho^{5}.$
We denote the subspace of elements of $\mathbb{C}\left[F_{r}\right]$ supported in the ball of radius $q$ by $\mathbb{C}_{\leq q}\left[F_{r}\right].$ Given a non–decreasing sequence of positive real numbers $\left(w(q)\right)_{q\in\mathbb{N}}$, we define the space
$\mathcal{S}_{w}\left(F_{r}\right)\overset{\mathrm{def}}{=}\Big{\{}x\in C^{*}\left(F_{r}\right)\ :\ x=\sum_{q}x_{q},\ x_{q}\in\mathbb{C}_{\leq q}\left[F_{r}\right],\ \sum_{q}w(q)\|x_{q}\|_{C^{*}\left(F_{r}\right)}<\infty\Big{\}}.$ (6)
If $w(q)=(1+q)^{i}$ for some $i$ then we write $\mathcal{S}_{i}\left(F_{r}\right)$ in place of $\mathcal{S}_{w}\left(F_{r}\right).$
$\mathcal{S}_{w}\left(F_{r}\right)$ is always a Banach space when equipped with the norm
$\|x\|_{w}\overset{\mathrm{def}}{=}\inf\Big{\{}\sum_{q}w(q)\|x_{q}\|_{C^{*}\left(F_{r}\right)}\ :\ x=\sum_{q}x_{q}\in\mathcal{S}_{w}\left(F_{r}\right)\Big{\}}.$
2.7.2 Temperedness and strong convergence
The following definition and proposition can be found in [MAgeeDelaSalle, Section 5]. Fix a generating set of size $r$ for the free group $F_{r}.$
Definition 2.19.
A function $u:F_{r}\to\mathbb{C}$ is called tempered if
$\limsup_{n\to\infty}|u\left(\left(x^{*}x\right)^{n}\right)|^{\frac{1}{2n}}\leq\|\lambda(x)\|_{\mathrm{op}}$
for every $x\in\mathbb{C}\left(F_{r}\right),$where $\lambda$ is the left regular representation.
The above property is a part of a criterion for strong convergence of random representations.
Proposition 2.20.
Let $u_{n}:F_{r}\to\mathbb{C}$ be functions, $\pi_{n}$ a sequence of random unitary representations of $F_{r}$ with finite and non–random dimension. Let $\epsilon_{n}>0.$ If the following conditions are satisfied:
  • $\lim_{n\to\infty}\epsilon_{n}=0,$
  • we have $|\mathbb{E}\mathrm{Tr}\left(\pi_{n}(x)\right)-u_{n}(x)|\leq\epsilon_{n}\exp\left(\frac{q}{\log\left(2+q\right)^{2}}\right)\|x\|_{C^{*}\left(F_{r}\right)}$ for every $q$ and every $x\in\mathbb{C}_{\leq q}\left[F_{r}\right],$ and
  • $u_{n}$ is tempered and there is a polynomial $P_{n}$ such that, for every q and every $x\in\mathbb{C}_{\leq q}\left[F_{r}\right],$ $|u_{n}(x)|\leq P_{n}(q)\|x\|_{C^{*}\left(F_{r}\right)}.$
Then, for every $y\in\mathbb{C}\left[F_{r}\right]$ and for every $\delta>0,$
$\mathbb{P}\left[\left\|\pi_{n}(y)\right\|>\left\|\lambda(y)\right\|+\delta\right]\leq C\left(y,\delta\right)\epsilon_{n}$
for some constant $C(y,\delta).$ In particular,
$\lim_{n}\mathbb{P}\left[\left\|\pi_{n}(y)\right\|>\left\|\lambda(y)\right\|+\delta\right]=0.$ (7)
To prove that a function is tempered, we can use the following proposition (stated here for free groups only), from [MAgeeDelaSalle, Proposition 6.3].
Proposition 2.21.
Let $u:F_{r}\to\mathbb{C}$ and assume that, for every reasonable probability measure $\mu$ on $F_{r}$, if $\left(g_{n}\right)_{n}$ is the associated random walk on $F_{r},$
$\limsup_{n\to\infty}\left(\mathbb{E}\left|u\left(g_{n}\right)\right|\right)^{\frac{1}{n}}\leq\rho(\mu).$
Then $u$ is tempered.
Remark 2.22.
That this proposition holds for free groups is a result of Haagerup’s inequality [Haagerup, Lemma 1.5], which asserts that free groups with their standard generating sets have the rapid decay property.
2.8 Schreier graphs
Definition 2.23.
Let $G$ be a finite group acting on a finite set $X$ and let $g_{1},\dots,g_{r}\in G.$ The Schreier graph $\mathrm{Sch}\left(S_{n}\curvearrowright X,g_{1},\dots,g_{r}\right)$ is the $2r$–regular graph on $|X|$ vertices, consisting of vertex set $X$ and, for every vertex $x\in X,$ for each $g_{i},$ we add an edge between $x$ and $g_{i}x.$ Note that we allow multiple edges and loops.
2.9 Word maps
Given a word $w\in F_{r}=\left\langle x_{1},\dots,x_{r}\right\rangle$ and a (finite) group $G,$ one obtains a word map
$w:G^{r}\to G$
by substitutions. For example, if $w=\left[x_{1},x_{2}\right]=x_{1}x_{2}x_{1}^{-1}x_{2}^{-1}\in F_{2}$ and $g,h\in G,$ then $w\left(g,h\right)=ghg^{-1}h^{-1}\in G.$
For any finite group $G$ and class function $\chi:G\to\mathbb{R},$ we define
$\mathbb{E}_{w}\left[\chi\right]\overset{\mathrm{def}}{=}\mathop{\mathbb{E}}_{g_{1},\dots,g_{m}\in G}\left[\chi\left(w\left(g_{1},\dots,g_{m}\right)\right)\right].$
In this paper, we will only be considering word maps for the symmetric group $S_{n}.$
Remark 2.24.
$\mathbb{E}_{w}\left[\chi\right]$ is equivalent to $\mathop{\mathbb{E}}_{\mathrm{Hom}\left(F_{r},S_{n}\right)}\left[\chi\left(\phi(w)\right)\right]$, where $\phi\in\mathrm{Hom}(F_{r},S_{n})$ is chosen uniformly randomly by choosing uniformly random i.i.d. $\sigma_{1},\dots,\sigma_{r}\in S_{n}$ and mapping each $x_{i}\mapsto\sigma_{i},$ as is described in §1.1.
3 Proof of Theorem 1.9 from Theorem 1.5
The results of this section are almost completely analogous to those of [MAgeeDelaSalle, Section 7]. Throughout, $C$ will denote a constant that does not depend on anything, but that may change from line to line. We require our bound on the expected irreducible stable character of a $w$–random permutation from Theorem 1.5, as well as our formulation of this expected character as a ratio of two polynomials in $\frac{1}{n},$ from Theorem 4.1.
Fix any integer $k>0$ and define
$\Sigma_{n,k}\overset{\mathrm{def}}{=}\bigoplus_{\lambda\vdash k}V^{\lambda^{+}(n)}.$
Then $\Sigma_{n,k}\subseteq\left(\mathbb{C}^{n}\right)^{\otimes k},$ so that
$\dim\left(\Sigma_{n,k}\right)\leq n^{k}.$ (8)
The following theorem follows immediately by combining Theorem 1.5 and Theorem 4.1.
Theorem 3.1.
For every word $w\in F_{r},$ there is a rational function $\phi_{w}\in\mathbb{Q}\left[x\right]$ such that
  1. 1.
    For $n\geq k\max\left(l(w),2\right),$
    $\phi_{w}\left(\frac{1}{n}\right)=\frac{1}{n^{k}}\mathbb{E}_{w}\left[\mathrm{Tr_{\Sigma_{n,k}}}\right].$
  2. 2.
    If $w$ is not the identity and $l(w)\leq q,$ then $g_{q,k}\phi_{w}$ is a polynomial of degree $\leq D_{q}=3kq+q^{2}.$
  3. 3.
    If $w$ is not a proper power then, for all $i<2k,$
    $\phi_{w}^{(i)}(0)=0.$
We collect some facts about the polynomial
$g_{L,k}(x)=\prod_{c=1}^{L}(1-cx)^{L}\left[\prod_{j=1}^{2k}\left(1-(j-1)x\right)\right]^{L}$
in the following lemma.
Lemma 3.2.
Fix an integer $L>0.$ Then, for every $t$ satisfying
$0\leq t\leq\frac{1}{2}\frac{1}{L^{2}(L+1)+L(2k-1)(2k)}$
and every integer $i\geq 0:$
  1. 1.
    $\frac{1}{2}\leq g_{L,k}(t)\leq 1$
  2. 2.
    $|g_{L,k}^{(i)}(t)|\leq(i+1)\left(L^{3}+L(2k-1)^{2}\right)^{i}$
  3. 3.
    $\left|\left(\frac{1}{g_{L,k}}\right)^{(i)}(t)\right|\leq(2i)!!2^{2i+1}\left(L^{3}+L(2k-1)^{2}\right)^{i}.$
Proof.
Fix any $0\leq t\leq\frac{1}{2}\frac{1}{L^{2}(L+1)+L(2k-1)(2k)}.$ That $g_{L,k}(t)\leq 1$ follows immediately from the fact that
$|t|<\min\Big{\{}\frac{1}{L},\frac{1}{2k-1}\Big{\}}.$
To see that $\frac{1}{2}\leq g_{L,k}(t),$ observe that, in the regime $0<z<\frac{1}{2},$ we have $(1-z)\geq e^{-2z}.$ Since
$t<\min\Big{\{}\frac{1}{2L},\frac{1}{2(2k-1)}\Big{\}},$
we have
$g_{L,k}(t)$
$\geq$ $\exp\left(-2\left(\sum_{c=1}^{L}Lct+\sum_{j=1}^{2k}L(j-1)t\right)\right)$
$=$ $\exp\left(-2Lt\left(1+\dots+L+1+\dots+2k-1\right)\right)$
$=$ $\exp\left(-2Lt\left(\frac{L(L+1)}{2}+\frac{(2k-1)2k}{2}\right)\right)$
$=$ $\exp\left(-t\left(L^{2}(L+1)+L(2k-1)(2k)\right)\right)$
$\geq$ $\frac{1}{2},$
where the final inequality follows from the fact that
$t\leq\frac{1}{2}\frac{1}{L^{2}(L+1)+L(2k-1)(2k)}<\frac{-\log\left(\frac{1}{2}\right)}{L^{2}(L+1)+L(2k-1)(2k)}.$
For Part 2, we begin by rewriting
$g_{L,k}(t)=\prod_{z=1}^{L^{2}+2kL}\left(1-b_{z}t\right)$
and then differentiate using the Leibniz rule. The $i^{\mathrm{th}}$ derivative is a sum of terms in which some terms are derived once (so are equal to $-b_{z}$) and some are not derived. Each factor that is not derived is of the form $1-b_{z}t,$ which belongs to $(0,1].$ So we can bound the derivative using only the terms that are derived:
$\left|g_{L,k}^{(i)}(t)\right|$
$\leq$ $\sum_{s=0}^{i}\sum_{\underset{\mathrm{distinct}}{u_{1},\dots,u_{s}}}\prod_{\alpha=1}^{s}b_{u_{\alpha}}$
$\leq$ $\sum_{s=0}^{i}\left(\sum_{z=1}^{L^{2}+2kL}b_{z}\right)^{s}$
$=$ $\sum_{s=0}^{i}\left(\sum_{c=1}^{L}Lc+\sum_{j=1}^{2k}L(j-1)\right)^{s}$
$\leq$ $\sum_{s=0}^{i}\left(L^{3}+L(2k-1)^{2}\right)^{s}$
$\leq$ $(i+1)\left(L^{3}+L(2k-1)^{2}\right)^{i}.$
To prove Part 3, we use that the $i^{\mathrm{th}}$ derivative of $\frac{1}{g_{L,k}}$ can be written as a product of $(2i)!!$! terms of the form
±$\frac{\pm g_{L,k}^{(\alpha_{1})}\dots g_{L,k}^{(\alpha_{i})}}{g_{L,k}^{i+1}},$
where $\alpha_{1},\dots,\alpha_{i}\in\mathbb{Z}_{\geq 0}$ and $\alpha_{1}+\dots+\alpha_{i}=i.$
By Part 2, each term is bounded by
$\frac{\prod_{j}\left(\alpha_{j}+1\right)\left(L^{3}+L(2k-1)^{2}\right)^{\alpha_{j}}}{g_{L,k}^{i+1}}$
and, by Part 1, this is bounded by
$2^{i+1}\prod_{j}\left(\alpha_{j}+1\right)\left(L^{3}+L(2k-1)^{2}\right)^{\alpha_{j}},$
which is then bounded by
$2^{2i+1}\left(L^{3}+L(2k-1)^{2}\right)^{i},$
from which Part 3 follows. ∎
Each $w\in F_{r}$ defines a map $w\mapsto\phi_{w}.$ We extend this by linearity to define a map $x\mapsto\phi_{x}$ for each
$x=\sum_{w\in F_{r}}x(w)w\in\mathbb{C}\left[F_{r}\right].$
Then we can prove the following.
Lemma 3.3.
For any $x\in\mathbb{C}_{\leq q}\left[F_{r}\right]$, for any $i\leq 2k$ and for $n\geq kq$,
$\sup_{t\in\left[0,\frac{1}{2D_{q}^{2}}\right]}\frac{\left|\phi_{x}^{(i)}(t)\right|}{i!}\leq h(i,q)||x||_{C^{*}\left(F_{r}\right)},$
where
$h(i,q)=4\left(\frac{CD_{q}^{4}}{i^{2}}\right)^{i}.$
Proof.
Let $P=g_{q,k}\phi_{x}.$ Then $P$ is a polynomial of bounded degree $d\leq D_{q}$ and so we can bound its derivatives using Lemma 2.16 if we can bound $P$ itself. For $n\geq kq,$ we have $0<g_{q,k}\left(\frac{1}{n}\right)<1,$ so that
$\left|P\left(\frac{1}{n}\right)\right|$
$\leq$ $\left|\phi_{x}\left(\frac{1}{n}\right)\right|$
$=$ $\frac{1}{n^{k}}\left|\mathbb{E}_{x}\left[\mathrm{Tr_{\Sigma_{n,k}}}\right]\right|$
$\underset{(\dagger)}{\leq}$ $\frac{1}{n^{k}}\dim\left(\Sigma_{n,k}\right)||x||_{C^{*}\left(F_{r}\right)}$
[…]
The inequality ()($\dagger$)( † ) follows since the map $w\mapsto\Sigma_{n,k}\left(w\left(\sigma_{1},\dots,\sigma_{r}\right)\right)$ is a unitary representation of $F_{r}$ for every $\left(\sigma_{1},\dots,\sigma_{r}\right)\in S_{n}^{r}.$ So we have
$\left|\left|\Sigma_{n,k}\left(x\left(\sigma_{1},\dots,\sigma_{r}\right)\right)\right|\right|\leq||x||_{C^{*}\left(F_{r}\right)}$
almost surely, from which we can bound
$\left|\mathbb{E}_{x}\left[\mathrm{Tr_{\Sigma_{n,k}}}\right]\right|\leq\dim\left(\Sigma_{n,k}\right)||x||_{C^{*}\left(F_{r}\right)}.$
By Lemma 2.16 (and using Stirling’s formula), for any integer $j,$ we can bound
$\sup_{t\in\left[0,\frac{1}{2D_{q}^{2}}\right]}\frac{\left|P^{(j)}(t)\right|}{j!}\leq 2\left(\frac{CD_{q}^{4}}{j^{2}}\right)^{j}\left\|x\right\|_{C^{*}\left(F_{r}\right)}.$ (9)
We have
$\frac{\phi_{x}^{(i)}}{i!}=\sum_{j=0}^{\min(i,D_{q})}\begin{pmatrix}i\\ j\end{pmatrix}\frac{j!}{i!}\frac{P^{(j)}}{j!}\left(\frac{1}{g_{q,k}}\right)^{(i-j)}.$
Note that $i\leq 2k\leq D_{q}=q^{2}+3kq,$ so that $\min(i,D_{q})=i.$ With the bound (9) and Lemma 3.2 Part 3, for $t\in\left[0,\frac{1}{2D_{q}^{2}}\right]$, each term (besides $j\neq i$ ) is smaller than $\frac{1}{i}2\left(\frac{CD_{q}^{4}}{i^{2}}\right)^{i},$ so that
$\frac{\left|\phi_{x}^{(i)}(t)\right|}{i!}\leq 2\left(\frac{CD_{q}^{4}}{i^{2}}\right)^{i}+\frac{i}{i}2\left(\frac{CD_{q}^{4}}{i^{2}}\right)^{i},$
which proves the lemma. ∎
For each integer $i\geq 0,$ we define a map
$\psi_{i}:x\in\mathbb{C}\left[F_{r}\right]\mapsto\frac{\phi_{x}^{(k+i)}(0)}{(k+i)!}\in\mathbb{C}.$
We want to show that $\psi_{i}$ is tempered and satisfies the polynomial bound property in Proposition 2.20.
Lemma 3.4.
For every integer $i$ with $0\leq i\leq k$, there is a polynomial $P$ of degree $14k+4i+$ 1 such that $\left|\psi_{i}(x)\right|\leq P_{n}(q)\|x\|_{C^{*}\left(F_{r}\right)}$ for every $q$ and for every $x\in\mathbb{C}_{\leq q}\left[F_{r}\right].$
Proof.
By Lemma 3.3, for each $q$ and for each $x\in\mathbb{C}_{\leq q}\left[F_{r}\right],$ we have
$\left|\frac{\phi_{x}^{(k+i)}(0)}{(k+i)!}\right|\leq h(k+i,q)\|x\|_{C^{*}\left(F_{r}\right)}.$ (10)
Moreover,
$\sup_{q\geq 1}\frac{h(k+i,q)}{(1+q)^{4k+4i+1}}<\infty,$
from which the lemma follows. ∎
Lemma 3.5.
For any $i<k,$ the function $\psi_{i}$ is tempered.
Proof.
We will use Proposition 2.21 by showing that, for any reasonable probability measure $\mu$ with associated random walk $\left(g_{n}\right)_{n},$
$\limsup_{n}\left(\mathbb{E}|\psi_{i}(g_{n})|\right)^{\frac{1}{n}}\leq\rho(\mu).$
So, let $\mu$ be a reasonable probability measure, $(g_{n})_{n}$ the associated random walk on $F_{r}.$ If $\mu$ is supported in $\mathbb{C}_{\leq q}\left[F_{r}\right],$ then $g_{n}\in\mathbb{C}_{\leq qn}\left[F_{r}\right].$ Then we have
$\mathbb{E}|\psi_{i}(g_{n})|$
$\leq$ $C_{i}(1+qn)^{4k+4i+1}\mathbb{P}\left(\psi_{i}(g_{n})\neq 0\right)$
$=$ $C_{i}(1+qn)^{4k+4i+1}\mathbb{P}\left(g_{n}\ \mathrm{is\ a\ proper\ power}\right)$
$\leq$ $C_{i}C_{\mu}(1+qn)^{4k+4i+1}n^{5}\rho(\mu)^{n}.$
The first inequality follows from Lemma 3.4. Indeed, $g_{n}\in\mathbb{C}_{\leq qn}\left[F_{r}\right],$ so there is some constant $C_{i}$ for which $|\psi(g_{n})|\leq C_{i}\left(1+qn\right)^{4k+4i+1}.$ The final inequality follows from Proposition 2.18. It follows that
$\limsup_{n}\left(\mathbb{E}|\psi_{i}(g_{n})|\right)^{\frac{1}{n}}\leq\rho(\mu),$
so by Proposition 2.21, $\psi_{i}$ is tempered. ∎
The next step is showing that the second condition in Proposition 2.20 holds. We will need the following lemma.
Lemma 3.6.
For any $\epsilon>0,$ we have $\sup_{q\geq 1}h(2k,q)\exp\left(-\frac{q}{\log(2+q)^{2}}\right)\leq\left(C^{2}k^{12+\epsilon}\right)^{k}.$
Proof.
Fix $\epsilon>0.$ Given any $b>1,$ there exists $a>0$ such that
$\log\left(2+q\right)^{2}\leq q^{1/b}$
for all $q>a$.
So, for $q>a,$ we have
$4\left(CD_{q}^{4}\right)^{2k}\exp\left(\frac{-q}{\log\left(2+q\right)^{2}}\right)\leq 4\left(C\left(q^{2}+3kq\right)^{4}\right)^{2k}\exp\left(-q^{\frac{b-1}{b}}\right)$ (11)
which is bounded above by
$\left(C^{2}k^{12+\epsilon}\right)^{k}$ (12)
for sufficiently large $b$ [^2] To see this, differentiate the RHS of (11) to see that the maximum is obtained around $q=Ck^{\frac{b}{b-1}}$and substitute this in to $h(2k,q)$.
For $q\leq a$, we obtain the bound
$\left(C^{\prime}k^{2}\right)^{2k}$
by substituting $q=a$ in to $h(2k,q)$ and ignoring the exponential term (since it is $\leq 1$). This is less than (12) (when $C$ is large enough) from which the lemma follows. ∎
Lemma 3.7.
Let $w(q)=\exp\left(\frac{q}{\log\left(2+q\right)^{2}}\right).$ Then, for every $q\geq 1,$ for every $n\geq kq$ and for every $x\in\mathbb{C}_{\leq q}\left[F_{r}\right],$ and any $\epsilon>0,$
$\mathop{\left|\mathbb{E}\left[\mathrm{Tr}\left(\Sigma_{n,k}\left(x(\sigma_{1},\dots,\sigma_{r})\right)-\tau(x)\mathrm{Id}\right)\right]-\sum_{i=0}^{k-1}\frac{\psi_{i}(x)}{n^{i}}\right|}$ (13)
$\leq$ $\frac{\left(C^{2}k^{12+\epsilon}\right)^{k}}{n^{k}}w(q)\|x\|_{C^{*}\left(F_{r}\right)}.$
Proof.
Since $n\geq kq,$ then by Part 1 of Theorem 3.1, the LHS of (13) is equal to
$n^{k}\left|\phi_{x}\left(\frac{1}{n}\right)-\sum_{i=0}^{2k-1}\frac{\phi_{x}^{(i)}(0)}{i!n^{i}}\right|.$
By Taylor’s inequality, this is less than or equal to
$\frac{n^{k}}{n^{2k}(2k)!}\left|\phi_{x}^{(2k)}\left(\frac{1}{n}\right)\right|.$
If we further assume that $n\geq 2D_{q}^{2},$ then by Lemma 3.3, this is bounded by
$\frac{h(2k,i)}{n^{k}}\|x\|_{C^{*}\left(F_{r}\right)}.$ (14)
If $n\leq 2D_{q}^{2}$, then we can still bound the left hand side of (13) by
$2\dim\left(\Sigma_{n,k}\right)\|x\|_{C^{*}\left(F_{r}\right)}+\sum_{i=0}^{k-1}\frac{\left|\psi_{i}(x)\right|}{n^{i}}$
by using the triangle inequality. By (8) and Lemma 3.3, this is bounded above by
$n^{k}\left[h(0,q)+\sum_{i=0}^{k-1}\frac{1}{n^{k+i}}h(k+i,q)\right]\|x\|_{C^{*}\left(F_{r}\right)}.$
This is less than (14) whenever the constant $C$ is large enough.
By Lemma 3.6, we have
$\sup_{q\geq 1}h(2k,q)\exp\left(-\frac{q}{\log(2+q)^{2}}\right)\leq\left(C^{2}k^{12+\epsilon}\right)^{k}$
and the lemma follows. ∎
Proof of Theorem 1.9 .
Fix $\alpha<\frac{1}{12},$ say $\alpha=\frac{1}{12}-\epsilon^{\prime},$ with $0<\epsilon^{\prime}\leq\frac{1}{12}.$ For each $n,$ and for any $k\leq n^{\alpha}$, let $\Pi_{n,k}$ be a random representation of $F_{r}$ given by
$\Pi_{n,k}(w)=\Sigma_{n,k}\left(w\left(\sigma_{1},\dots,\sigma_{r}\right)\right).$
Then, by Lemma 3.7, for every $q$ and every $x\in\mathbb{C}_{\leq q}\left[F_{r}\right],$ we have
$\left|\mathbb{E}\mathrm{Tr}\left(\Pi_{n,k}(x)\right)-T_{n}(x)\right|\leq\epsilon_{n}w(q)\|x\|_{C^{*}\left(F_{r}\right)},$
where, for some $\epsilon$ satisfying $0<\epsilon<\frac{12\epsilon^{\prime}}{\frac{1}{12}-\epsilon^{\prime}}$, we have
$\epsilon_{n}=\frac{\left(C^{2}k^{12+\epsilon}\right)^{k}}{n^{k}}\leq\left(\frac{C^{2}}{n^{1-(12+\epsilon)\alpha}}\right)^{k}$
and
$T_{n}(x)=\dim\left(\Sigma_{n,k}\right)\tau(x)+\sum_{i=0}^{k-1}\frac{\psi_{i}(x)}{n^{i}}.$
By Lemma 3.5, $T_{n}$ is tempered, since it is the finite sum of tempered functions. Additionally, $T_{n}$ satisfies the polynomial bound
$\left|T_{n}(x)\right|\leq P_{n}(q)\|x\|_{C^{*}\left(F_{r}\right)}.$
So by Proposition 2.20, with $T_{n}$ in place of $u_{n}$, $\forall\delta>0$, and any $z\in\mathbb{C}\left[F_{r}\right],$
$\mathbb{P}\left[\left\|\Pi_{n,k}(z)\right\|>\left\|z\right\|_{C_{\lambda}^{*}(F_{r})}+\delta\right]\leq C(z,\delta)\left(\frac{C^{2}}{n^{1-(12+\epsilon)\alpha}}\right)^{k}.$
For $k_{n}\leq n^{\alpha},$ we have
$\mathbb{P}\left[\left\|\pi_{n,k_{n}}(z)\right\|>\left\|z\right\|_{C_{\lambda}^{*}(F_{r})}+\delta\right]$ (15)
$\leq$ $\sum_{1\leq k\leq n^{\alpha}}\mathbb{P}\left[\left\|\Pi_{n,k}(z)\right\|>\left\|z\right\|_{C_{\lambda}^{*}(F_{r})}+\delta\right].$
Write $A=1-(12+\epsilon)\alpha,$ which is $>0$ by our choice of $\epsilon.$ Then the RHS of (15) is equal to
$C(x,\delta)\left(\frac{C^{2}}{n^{A}}\right)\left(\frac{1-\left(\frac{C^{2}}{n^{A}}\right)^{n^{\alpha}-1}}{1-\left(\frac{C^{2}}{n^{A}}\right)}\right),$
which $\to 0$ as $n\to\infty.$ So, for any $k_{n}\leq n^{\alpha},$
$\mathbb{P}\left[\left\|\pi_{n,k_{n}}(z)\right\|>\left\|z\right\|_{C_{\lambda}^{*}(F_{r})}+\delta\right]\overset{n\to\infty}{\longrightarrow}0.$ (16)
By [MAgeeDelaSalle, Lemma 5.14], there exist $y_{1},\dots,y_{m}\in\mathbb{C}\left[F_{r}\right]$ and $\delta^{\prime}=\delta^{\prime}(\delta),$ such that if, for every $i,$
$\left\|\pi_{n,k_{n}}\left(y_{i}\right)\right\|\leq\left\|y_{i}\right\|_{C_{\lambda}^{*}(F_{r})}+\min(\delta,\delta^{\prime}),$ (17)
then
$\left\|\pi_{n,k_{n}}\left(z\right)\right\|\geq\left\|z\right\|_{C_{\lambda}^{*}(F_{r})}-\delta.$ (18)
By (16), we know that (17) holds with probability $\to 1$ as $n\to\infty,$ so then (18) also holds with probability $\to 1$ as $n\to\infty.$ Therefore,
$\mathbb{P}\left[\sup_{k_{n}\leq n^{\alpha}}\left|\left\|\pi_{n,k_{n}}\left(z\right)\right\|-\left\|z\right\|_{C_{\lambda}^{*}(F_{r})}\right|>\delta\right]\overset{n\to\infty}{\longrightarrow}0.$
4 Proof of Theorem 1.5
4.1 Overview of proof
Rather than working directly with the character $\chi^{\lambda^{+}(n)},$ we use Corollary 2.11 so that instead our task is to compute the bitrace of $\mathcal{Q}_{\lambda,n}\circ w\left(g_{1},\dots,g_{r}\right)$ on $\left(\mathbb{C}^{n}\right)^{\otimes k}.$
This amounts to counting the number of fixed points in $\left(\mathbb{C}^{n}\right)^{\otimes k}$ of this map. Recall that $\mathcal{Q}_{\lambda,n}^{2}=\mathcal{Q}_{\lambda,n}$ and that, since it is a linear combination of $p_{\pi},$ the action of $\mathcal{Q}_{\lambda,n}$ commutes with the action of any $g\in S_{n}$ on $\left(\mathbb{C}^{n}\right)^{\otimes k}.$ With this in mind, if $w=x_{i_{1}}^{\pm 1}\dots x_{i_{l(w)}}^{\pm 1}$, what we are actually interested in is
$\mathrm{btr}_{\left(\mathbb{C}^{n}\right)^{\otimes k}}\left(g_{i_{1}}^{\pm 1}\circ\mathcal{Q}_{\lambda,n}\circ\dots\circ g_{i_{l(w)}}^{\pm 1}\circ\mathcal{Q}_{\lambda,n}\right),$
where each $g_{i}\in S_{n}$ is a permutation substituted in for the generator $x_{i}.$
We compute the expected trace by using the Weingarten calculus for $S_{n}.$ A key component is our refinement of the usual Weingarten calculus for $S_{n},$ which uses the fact that we are operating within $A_{k}(n)$ to show that the contribution to the trace from all but a specific family of partitions is zero.
We obtain a combinatorial formula for the trace in §4.3. In §4.4, we construct graphs from the combinatorial data formula for the expected character. From these graphs we construct new graphs in which the asymptotic bound for our trace formula is encoded by the Euler characteristic.
In §4.5, we prove a variant of a theorem of Louder and Wilton that relates the Euler characteristic of a graph with the number of immersed $w$–cycles (see [LouderWilton, Theorem 1.2]) to obtain our final bound for the expected character.
4.2 Expected character as a ratio of polynomials in $\frac{1}{n}$
In addition to our main theorem on word maps, we give another formulation for the expected stable irreducible character of a $w$–random permutation as the ratio of two polynomials in $\frac{1}{n}.$ This formulation is required for the methods in §3.
For any $L,k\in\mathbb{Z}_{>0},$ we define
$g_{L,k}(x)\overset{\mathrm{def}}{=}\prod_{c=1}^{L}(1-cx)^{L}\left[\prod_{j=1}^{2k}\left(1-(j-1)x\right)\right]^{L}.$
Theorem 4.1.
Let $w\in F_{r}=\left\langle x_{1},\dots,x_{r}\right\rangle$ be a word in the free group with $r$ generators, $w\neq e$. Let $k\in\mathbb{Z}_{>0}$ and $\lambda\vdash k$. If $l(w)\leq q,$ there is a polynomial $P_{w,\lambda,q}\in\mathbb{Q}\left[x\right]$ such that, for $n\geq l(w)k,$
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=\frac{P_{w,\lambda,q}\left(\frac{1}{n}\right)}{g_{q,k}\left(\frac{1}{n}\right)},$
with $\deg\left(P_{w,\lambda,q}\right)\leq 3kq+q^{2}.$
The proof of Theorem 4.1 is given in §4.6 and it can be followed from (24).
4.3 Combinatorial integration
Fix a word $w\in F_{r}=\left\langle x_{1},\dots,x_{r}\right\rangle.$ We will assume that $w$ is not the identity and that $w$ is cyclically reduced. We will also assume that every generator appears in $w$. If this were not the case, then we can relabel the $x_{i}$ that do appear in $w$ and write $w\in F_{m}=\left\langle x_{i_{1}},\dots,x_{i_{m}}\right\rangle$ for some $m<r$ and then apply the exact same proof.
We will write
$w=f_{1}^{\epsilon_{1}}f_{2}^{\epsilon_{2}}\dots f_{l(w)}^{\epsilon(w)},$
where each $f_{i}\in\{x_{1},\dots,x_{r}\}$ and each $\epsilon_{i}\in\{1,-1\}.$ We will write $|w|_{x_{i}}$ for the number of $j$ such that $f_{j}=x_{i}.$
Suppose that $\{v_{p}\}$ is an orthonormal basis for $\mathcal{U}_{\lambda,n}\cong V^{\lambda^{+}(n)}\otimes V^{\lambda}.$ Then, for any $\left(g_{x_{1}},\dots,g_{x_{r}}\right)\in S_{n}^{r},$ we have
$\chi^{\lambda^{+}(n)}\left(w\left(g_{x_{1}},\dots,g_{x_{r}}\right)\right)=d_{\lambda}\mathrm{tr}_{\mathcal{U}_{\lambda,n}}\left(w\left(g_{x_{1}},\dots,g_{x_{r}}\right)\right),$
so that
$\chi^{\lambda^{+}(n)}\left(w\left(g_{x_{1}},\dots,g_{x_{r}}\right)\right)$ (19)
$=$ $d_{\lambda}\sum_{p_{i}}\left\langle g_{f_{1}}^{\epsilon_{1}}v_{p_{2}},\ v_{p_{1}}\right\rangle\left\langle g_{f_{2}}^{\epsilon_{2}}v_{p_{3}},\ v_{p_{2}}\right\rangle\dots\left\langle g_{f_{l(w)}}^{\epsilon_{l(w)}}v_{p_{1}},\ v_{p_{l(w)}}\right\rangle.$
For each $p$, write $v_{p}$ in the standard orthonormal basis of $\left(\mathbb{C}^{n}\right)^{\otimes k}$:
$v_{p}=\sum_{I}\beta_{p,I}e_{I}.$
Recall that $\mathcal{U}_{\lambda,n}\subset D_{k}(n),$ so that we may assume that the above sum is over all $I=\left(i_{1},\dots,i_{k}\right)$ with all indices distinct. This is an important observation.
We introduce some new notation to avoid the cumbersome general expression for $\mathop{\mathbb{E}}_{w}\left[\chi^{\lambda^{+}(n)}\right].$ For each $f\in\{x_{1},\dots,x_{r}\},$ we will write
$\sum_{I^{f}}$
in place of
$\sum_{I_{f}^{1},\dots I_{f}^{|w|_{f}}}.$
For any pair $I,J$ of multi–indices and $\epsilon\in\{1,-1\},$ we define
$(I/J)(\epsilon)=\begin{cases}J&\mathrm{if}\ \epsilon=1\\ I&\mathrm{if}\ \epsilon=-1.\end{cases}$
In our expression, we will simply write $(I/J)$ in place of $(I/J)(\epsilon)$ where it is clear which epsilon we are considering. To be even more clear, in the sum below, for each $f_{i},$ if $\epsilon_{i}=1,$ then the corresponding inner product has $e_{I_{f_{i}}^{z}}$ on the LHS, where $z\in\{1,\dots,|w|_{f_{i}}\}$ denotes the number of times $f_{i}$ has appeared in the subword $f_{1}^{\epsilon_{1}}\dots f_{i}^{\epsilon_{i}}.$ The corresponding pair of $\beta$–terms has $\beta_{p,I},$ with $\beta_{p,J}$ conjugated. If $\epsilon_{i}=-1,$ we swap the positions of $I$ and $J$ [^3] We use this notation so that, in the graph construction detailed in Section 4.4, $\left(J_{f}^{i}\right)_{j}$ always represents the initial vertex of an $f$–edge and $\left(I_{f}^{i}\right)_{j}$ always represents the terminal vertex.
We will write
$\prod_{f,w}\beta$
$=$ $\left(\beta_{p_{2},\left(I_{f_{1}}^{1}/J_{f_{1}}^{1}\right)}\right)\left(\bar{\beta}_{p_{1},\left(J_{f_{1}}^{1}/I_{f_{1}}^{1}\right)}\right)\dots$
$\dots\left(\beta_{p_{1},\left(I_{f_{l(w)}}^{|w|_{f_{l(w)}}}/J_{f_{l(w)}}^{|w|_{f_{l(w)}}}\right)}\right)\left(\bar{\beta}_{p_{l(w)},\left(J_{f_{l(w)}}^{|w|_{f_{l(w)}}}/I_{f_{l(w)}}^{|w|_{f_{l(w)}}}\right)}\right).$
With this notation, (19) is equal to
$d_{\lambda}\sum_{p_{i}}\sum_{I^{f},J^{f}}\left(\prod_{f,w}\beta\right)\left\langle g_{f_{1}}^{\epsilon_{1}}e_{\left(I_{f_{1}}^{1}/J_{f_{1}}^{1}\right)},\ e_{\left(J_{f_{1}}^{1}/I_{f_{1}}^{1}\right)}\right\rangle\dots$ (20)
$\dots\left\langle g_{f_{l(w)}}^{\epsilon_{l(w)}}e_{\left(I_{f_{l(w)}}^{|w|_{f_{l(w)}}}/J_{f_{l(w)}}^{|w|_{f_{l(w)}}}\right)},e_{\left(J_{f_{l(w)}}^{|w|_{f_{l(w)}}}/I_{f_{l(w)}}^{|w|_{f_{l(w)}}}\right)}\right\rangle.$
See the example below. Without losing generality, we will always assume that $f_{1}=x_{1}$ and that $\epsilon_{1}=1$.
Example 4.2.
Suppose $w=x_{1}x_{2}x_{1}^{-1}x_{2}^{-1}.$ Then $\chi^{\lambda^{+}(n)}\left(w\left(g_{x_{1}},g_{x_{2}}\right)\right)$ is equal to
$d_{\lambda}\sum_{p_{i}}\sum_{\begin{aligned} I_{x_{1}}^{1},I_{x_{1}}^{2},I_{x_{2}}^{1},I_{x_{2}}^{2},\\ J_{x_{1}}^{1},J_{x_{1}}^{2},J_{x_{2}}^{1},J_{x_{2}}^{2}\end{aligned}}\left(\beta_{p_{2},I_{x_{1}}^{1}}\right)\left(\bar{\beta}_{p_{1},J_{x_{1}}^{1}}\right)\left(\beta_{p_{3},I_{x_{2}}^{1}}\right)\left(\bar{\beta}_{p_{2},J_{x_{2}}^{1}}\right)\left(\beta_{p_{4},J_{x_{1}}^{2}}\right)\left(\bar{\beta}_{p_{3},I_{x_{1}}^{2}}\right)$
$\left(\beta_{p_{1},J_{x_{2}}^{2}}\right)\left(\bar{\beta}_{p_{4},I_{x_{2}}^{2}}\right)\left\langle g_{x_{1}}e_{J_{x_{1}}^{1}},e_{I_{x_{1}}^{1}}\right\rangle\left\langle g_{x_{2}}e_{J_{x_{2}}^{1}},e_{I_{x_{2}}^{1}}\right\rangle\left\langle g_{x_{1}}^{-1}e_{I_{x_{1}}^{2}},e_{J_{x_{1}}^{2}}\right\rangle\left\langle g_{x_{2}}^{-1}e_{I_{x_{2}}^{2}},e_{J_{x_{2}}^{2}}\right\rangle.$
We rewrite the inner product terms as products of matrix coefficients – for example,
$\left\langle g_{x_{1}}e_{J_{x_{1}}^{1}},e_{I_{x_{1}}^{1}}\right\rangle=\left(g_{x_{1}}\right)_{\left(I_{x_{1}}^{1}\right)_{1}\left(J_{x_{1}}^{1}\right)_{1}}.\dots\left(g_{x_{1}}\right)_{\left(I_{x_{1}}^{1}\right)_{k}\left(J_{x_{1}}^{1}\right)_{k}}.$
Rearranging and taking the expectation over $g_{x_{1}},\dots,g_{x_{r}}\in S_{n},$ we obtain from (20) that $\mathop{\mathbb{E}}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ is equal to
$d_{\lambda}\sum_{p_{i}}\sum_{I^{f},J^{f}}\left(\prod_{f,w}\beta\right)\prod_{f\in\{x_{1},\dots,x_{r}\}}\int_{g_{I^{f},J^{f}}},$ (21)
where
$\int_{g_{I^{f},J^{f}}}\overset{\mathrm{def}}{=}\int_{S_{n}}\prod_{i=1}^{|w|_{f}}\left(g_{f}\right)_{\left(I_{f}^{i}\right)_{1}\left(J_{f}^{i}\right)_{1}}.\dots\left(g_{f}\right)_{\left(I_{f}^{i}\right)_{k}\left(J_{f}^{i}\right)_{k}}dg_{f}.$
For each $f\in\{x_{1},\dots,x_{r}\}$ and for each fixed collection of multi–indices $I_{f}^{1},\dots,I_{f}^{|w|_{f}},J_{f}^{1},\dots,J_{f}^{|w|_{f}}$, this integral can be computed using the Weingarten calculus for the symmetric group. Using Theorem 2.14 with $m=|w|_{f}k,$ we have
$\int_{S_{n}}\prod_{i=1}^{|w|_{f}}\left(g_{f}\right)_{\left(I_{f}^{i}\right)_{1}\left(J_{f}^{i}\right)_{1}}.\dots\left(g_{f}\right)_{\left(I_{f}^{i}\right)_{k}\left(J_{f}^{i}\right)_{k}}dg_{f}$
$=$ $\sum_{\sigma_{f},\tau_{f}\in\mathrm{Part}\left(\left[|w|_{f}k\right]\right)}\delta_{\sigma_{f}}\left(J_{f}^{1}\sqcup J_{f}^{2}\sqcup\dots\sqcup J_{f}^{|w|_{f}}\right)\delta_{\tau_{f}}\left(I_{f}^{1}\sqcup I_{f}^{2}\sqcup\dots\sqcup I_{f}^{|w|_{f}}\right)$
$\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right).$
Now we give an improvement over the usual Weingarten calculus that simplifies the above equation greatly. The benefit of the improvement is that, instead of summing over all set partitions $\sigma_{f},\tau_{f}\in\mathrm{Part}\left(\left[|w|_{f}k\right]\right),$ we show that we need only sum over set partitions $\sigma_{f},\tau_{f}$ that have a specific structure, since the contribution to (21) from the set partitions without this structure is zero.
To each partition we associate a diagram (in a similar way to §2.2) – the diagram has $|w|_{f}$ rows, each containing exactly $k$ vertices. The vertices in row $R$ are labeled from $(R-1)k+1$ to $Rk$, and two vertices are connected if and only if the corresponding vertex labels are in the same block of of the partition.
We will also refer to the diagrams corresponding to $\sigma_{f}$ and $\tau_{f}$ as $\sigma_{f}$ and $\tau_{f}$ where this does not cause confusion.
Lemma 4.3.
For each $f\in\{x_{1},\dots,x_{r}\},$ if $\sigma_{f}$ has any two vertices from the same row connected, then
$\delta_{\sigma_{f}}\left(J_{f}^{1}\sqcup J_{f}^{2}\sqcup\dots\sqcup J_{f}^{|w|_{f}}\right)=0.$
The same is true for $\tau_{f}$ and $I_{f}^{1}\sqcup I_{f}^{2}\sqcup\dots\sqcup I_{f}^{|w|_{f}}.$
Proof.
Since $e_{J_{f}^{1}}\in D_{k}(n)=\left\langle e_{i_{1}}\otimes\dots\otimes e_{i_{k}}:|\{i_{1},\dots,i_{k}\}|=k\right\rangle$, if any of the vertices in the top row of $\sigma_{f}$ are in the same block, then
$\delta_{\sigma_{f}}\left(J_{f}^{1}\sqcup J_{f}^{2}\sqcup\dots\sqcup J_{f}^{|w|_{f}}\right)=0.$
Repeating this argument for $J_{f}^{2},\dots J_{f}^{|w|_{f}}$ proves the claim for $\sigma_{f}$ and then repeating it for $I_{f}^{1}\sqcup I_{f}^{2}\sqcup\dots\sqcup I_{f}^{|w|_{f}}$ proves the claim for $\tau_{f}.$
The next lemma asserts that every vertex must be connected to at least one other vertex
Lemma 4.4.
For any $f\in\{x_{1},\dots,x_{r}\},$ if $\sigma_{f}$ has any singletons, then
$d_{\lambda}\sum_{p_{i}}\sum_{I^{f},J^{f}}\left(\prod_{f,w}\beta\right)$
$\Bigg{[}\sum_{\tau_{f}\in\mathrm{Part}\left(\left[|w|_{f}k\right]\right)}\delta_{\sigma_{f}}\left(J_{f}^{1}\sqcup J_{f}^{2}\sqcup\dots\sqcup J_{f}^{|w|_{f}}\right)\delta_{\tau_{f}}\left(I_{f}^{1}\sqcup I_{f}^{2}\sqcup\dots\sqcup I_{f}^{|w|_{f}}\right)$
$\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\Bigg{]}=0.$
The same is true when we swap $\sigma_{f}$ and $\tau_{f}.$
Proof.
Without loss of generality, suppose that the vertex $q$ in the first row of vertices of $\sigma_{f}$ is a singleton. For all variables in the sum fixed except for $J_{f}^{1},J_{f}^{2},\dots,J_{f}^{|w|_{f}},$ assuming
$\delta_{\tau_{f}}\left(I_{f}^{1}\sqcup I_{f}^{2}\sqcup\dots\sqcup I_{f}^{|w|_{f}}\right)\neq 0,$
then the sum is equal to
$\sum_{J_{f}^{1},\dots,J_{f}^{|w|_{f}}}\left(\bar{\beta}_{p_{1},J_{f}^{1}}\right)\dots\left(\beta_{p,J_{f}^{|w|_{f}}}\right)\delta_{\sigma_{f}}\left(J_{f}^{1}\sqcup J_{f}^{2}\sqcup\dots\sqcup J_{f}^{|w|_{f}}\right),$
multiplied by some constant coming from the other (fixed) $\beta$–terms and the Weingarten function $\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right).$
Then, for every fixed index $\left(J_{f}^{1}\right)_{1},\dots,\left(J_{f}^{1}\right)_{q-1},\left(J_{f}^{1}\right)_{q+1},\dots,\left(J_{f}^{1}\right)_{k}$ and every fixed multi–index $J_{f}^{2},\dots,J_{f}^{|w|_{f}}$ satisfying $\sigma_{f},$ this is (some constant multiplied by)
$\sum_{\left(J_{f}^{1}\right)_{q}=1}^{n}\bar{\beta}_{p_{1},J_{f}^{1}}.$
This is exactly the conjugate of the coefficient of $e_{(J_{f}^{1})_{1}}\otimes\dots\otimes e_{(J_{f}^{1})_{q-1}}\otimes e_{(J_{f}^{1})_{q+1}}\otimes\dots e_{(J_{f}^{1})_{k}}$ in $T_{q}\left(v_{p_{1}}\right).$ But then, $v_{p_{1}}$ belongs to an orthonormal basis for $\mathcal{U}_{\lambda,n},$ and $\mathcal{U}_{\lambda,n}\subset A_{k}(n).$ In particular, $v_{p}\in\mathrm{Ker}\left(T_{q}\right),$ so that
$\sum_{\left(J_{f}^{1}\right)_{q}=1}^{n}\bar{\beta}_{p_{1},J_{f}^{1}}=0.$
Henceforth, we will denote by $\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right)$ the set of set partitions of $\left[|w|_{f}k\right]$ for which the corresponding diagram has no singletons and has no two vertices in the same row in the same connected component.
We will write
$\sum_{\sigma_{f},\tau_{f}}^{\star}$
to indicate that the sum is over $\sigma_{f},\tau_{f}\in\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right).$
Remark 4.5.
It follows that if any of $x_{1},\dots,x_{r}$ appear only once in $w$, then there are no such set partitions in $\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right)$, and so in these cases $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=0.$ Such words are primitive and we know from [HananyPuder] that the expected stable irreducible character for primitive words is zero, unless $k=0.$ From now on, we will only be considering non–primitive words.
For any given $\sigma_{f}\in\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right)$ (and similarly for $\tau_{f}$), and any collection of multi–indices $J_{f}^{1},\dots,J_{f}^{|w|_{f}},$ we will write
$\sigma_{f}\leftrightarrow J_{f}^{\Sigma}$
to indicate that
$\delta_{\sigma_{f}}\left(J_{f}^{1}\sqcup J_{f}^{2}\sqcup\dots\sqcup J_{f}^{|w|_{f}}\right)=1.$
Therefore, our new expression for $\mathop{\mathbb{E}}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ is
$d_{\lambda}\sum_{\sigma_{f},\tau_{f}}^{\star}\left(\prod_{f\in\{x_{1},\dots,x_{r}\}}\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\right)\sum_{p_{1}}\sum_{\begin{aligned} \sigma_{f}\leftrightarrow J_{f}^{\Sigma},\\ \tau_{f}\leftrightarrow I_{f}^{\Sigma}\end{aligned}}\left(\prod_{f,w}\beta\right).$ (22)
This expression can be further simplified using the projection $\mathcal{Q}_{\lambda,n}.$ Consider Example 4.2. For all terms fixed in the expression except for, say, $p_{2},$ the sum that follows is
$\sum_{p_{2}}\left(\beta_{p_{2},I_{x_{1}}^{1}}\right)\left(\bar{\beta}_{p_{2},J_{x_{2}}^{1}}\right)$
$=$ $\sum_{p_{2}}\left\langle e_{I_{x_{1}}^{1}},v_{p_{2}}\right\rangle\left\langle v_{p_{2}},e_{J_{x_{2}}^{1}}\right\rangle$
$=$ $\left\langle\mathcal{Q}_{\lambda,n}\left(e_{I_{x_{1}}^{1}}\right),e_{J_{x_{2}}^{1}}\right\rangle$
$=$ $\sum_{\pi\in\mathrm{Part}\left([2k]\right)}c(n,k,\lambda,\pi)\left\langle p_{\pi}\left(e_{I_{x_{1}}^{1}}\right),e_{J_{x_{2}}^{1}}\right\rangle.$
Repeating this and computing the sums over each $p_{i},$ the expression in Example 4.2 becomes
$d_{\lambda}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{w}}\left(\prod_{i=1}^{4}c(n,k,\lambda,\pi_{i})\right)\left(\prod_{f\in\{x_{1},x_{2}\}}\mathrm{Wg}_{n,(|w|_{f}k)}\left(\sigma_{f},\tau_{f}\right)\right)$ (23)
$\sum_{\sigma_{f}\leftrightarrow J_{f}^{\Sigma},\ \tau_{f}\leftrightarrow I_{f}^{\Sigma}}\left\langle p_{\pi_{1}}\left(e_{I_{x_{1}}^{1}}\right),e_{J_{x_{2}}^{1}}\right\rangle\left\langle p_{\pi_{2}}\left(e_{I_{x_{2}}^{1}}\right),e_{I_{x_{1}}^{2}}\right\rangle$
$\left\langle p_{\pi_{3}}\left(e_{J_{x_{1}}^{2}}\right),e_{I_{x_{2}}^{2}}\right\rangle\left\langle p_{\pi_{4}}\left(e_{J_{x_{2}}^{2}}\right),e_{J_{x_{1}}^{1}}\right\rangle.$
Now, recall that the multi–indices in our sum must have all indices distinct. Therefore, if $\pi_{1}$ has any two vertices in the top row connected, or any two vertices in the bottom row connected, then the term
$\left\langle p_{\pi_{1}}\left(e_{I_{x_{1}}^{1}}\right),e_{J_{x_{2}}^{1}}\right\rangle=0$
for any pair of multi–indices. Similarly for $\pi_{2},\ \pi_{3}$ and $\pi_{4}.$ So the only summands that contribute to (23) are those in which each $\pi_{i}\leq\hat{\pi}_{i},$ where $\hat{\pi}_{i}=\iota(\rho)$ for some permutation $\rho\in S_{k}.$ In this case, we can replace the sum over the inner product terms in (23) by the number of multi–indices $I_{x_{1}}^{1},I_{x_{1}}^{2},I_{x_{2}}^{1},I_{x_{2}}^{2},J_{x_{1}}^{1},J_{x_{1}}^{2},J_{x_{2}}^{1},J_{x_{2}}^{2}$ that simultaneously satisfy $\sigma_{f}\leftrightarrow J_{f}^{\Sigma},\ \tau_{f}\leftrightarrow I_{f}^{\Sigma}$ as well as all of the inner product terms being equal to one.
We write
$\mathrm{Part}_{\leq S_{k}}\left([2k]\right)\overset{\mathrm{def}}{=}\big{\{}\pi\in\mathrm{Part}\left([2k]\right):\ \pi\leq\iota(\rho),\ \rho\in S_{k}\big{\}}.$
Additionally, we will write
$\sum_{\pi}^{\leq S_{k}}$
to indicate that we are summing over $\pi\in\mathrm{Part}_{\leq S_{k}}\left([2k]\right).$
This argument applies for any non–identity, cyclically reduced, non–primitive word $w.$ With this notation and the definition below, this proves Theorem 4.7.
Definition 4.6.
Given, for each $f\in\{x_{1},\dots,x_{r}\}$, a collection of set partitions $\sigma_{f},\tau_{f}\in\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right)$ and a collection of set partitions $\pi_{1},\dots,\pi_{l(w)}\in\mathrm{Part}_{\leq S_{k}}\left([2k]\right),$ we write
$\mathcal{N}\left(\sigma_{x_{1}},\tau_{x_{1}},\dots,\sigma_{x_{r}},\tau_{x_{r}},\pi_{1},\dots,\pi_{l(w)}\right)=\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right),$
for the number of multi–indices $I_{f}^{1},\dots I_{f}^{|w|_{f}},J_{f}^{1},\dots,J_{f}^{|w|_{f}},$ with all indices distinct, satisfying:
  • $\sigma_{f}\leftrightarrow J_{f}^{\Sigma}$,
  • $\tau_{f}\leftrightarrow I_{f}^{\Sigma}$,
  • $\left\langle p_{\pi_{1}}\left(e_{I_{x_{1}}^{1}}\right),\ e_{(I/J)_{f_{2}}}\right\rangle=1$,
\ \ \ \ \ $\vdots$
  • $\left\langle p_{\pi_{l(w)}}\left(e_{(I/J)_{f_{l(w)}}}\right),\ e_{J_{x_{1}}^{1}}\right\rangle=1,$
where $(I/J)$ is used to denote the correct multi–index arising from the $\beta$–terms.
Theorem 4.7.
Suppose $k\in\mathbb{Z}_{>0}.$ For any $\lambda\vdash k$ and every cyclically reduced, non–identity, non–primitive word $w\in F_{r},$
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ (24)
$=$ $d_{\lambda}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{i=1}^{l(w)}c(n,k,\lambda,\pi_{i})\right)\left(\prod_{f\in\{x_{1},\dots,x_{r}\}}\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\right)$
$\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right).$
Combining (24) with Remark 2.12 and Lemma 2.15, the following bound is immediate.
Corollary 4.8.
Suppose $k\in\mathbb{Z}_{>0}.$ For any $\lambda\vdash k$ and every cyclically reduced, non–identity, non–primitive word $w\in F_{r},$
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]\ll_{k,l(w)}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}n^{-\sum_{i=1}^{l(w)}\mathrm{del}\left(\pi_{i}\right)}n^{-\sum_{f}|\sigma_{f}\wedge\tau_{f}|}\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right).$
4.4 Graphical Interpretation
We construct a graph for each collection of set partitions $\sigma_{f},\tau_{f},\pi_{1},\dots,\pi_{l(w)}.$ This is similar to the construction of a surface from a matching datum by Magee in [Magee2021]. Indeed, the graph is essentially the same as the $1$–skeleton of the surfaces constructed there.
The graph we construct will be denoted $G\left(\sigma_{f},\tau_{f},\pi_{i}\right).$ The information in (24) will be encoded in the properties of $G\left(\sigma_{f},\tau_{f},\pi_{i}\right),$ allowing us to simplify further by analyzing the graph rather than dealing with tricky combinatorial arguments. Ultimately, from $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ we will construct a new graph from which we derive Theorem 1.5.
$G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ is a graph with $2kl(w)$ vertices, separated in to $k$ distinct subsets, each containing exactly $2l(w)$ vertices. We number the subsets from $1$ to $k$. For each $i\in\{1,\dots,k\}$, the vertices in subset $i$ are labeled
$\left(I_{f}^{1}\right)_{i},\dots\left(I_{f}^{|w|_{f}}\right)_{i},\left(J_{f}^{1}\right)_{i},\dots,\left(J_{f}^{|w|_{f}}\right)_{i}$ for each $f\in\{x_{1},\dots,x_{r}\}$.
Within each subset $i$, for each $f\in\{x_{1},\dots,x_{r}\}$ and for each $j\in\{1,\dots,|w|_{f}\},$ we draw a directed, $f$–labeled edge from the vertex labeled $\left(J_{f}^{j}\right)_{i}$ to the vertex labeled $\left(I_{f}^{j}\right)_{i}$. This gives a total of $kl(w)$ directed edges. We refer to the $i^{\mathrm{th}}$ subset of vertices as the $i^{\mathrm{th}}$ $w$–loop, see Figure 1.
$\scriptstyle\left(J_{x_{1}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{1}\right)_{1}$$\scriptstyle\left(J_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(J_{x_{2}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{2}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{3}\right)_{1}$$\scriptstyle\left(J_{x_{1}}^{3}\right)_{1}$$\scriptstyle\left(I_{x_{2}}^{2}\right)_{1}$$\scriptstyle\left(J_{x_{2}}^{2}\right)_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle\left(J_{x_{1}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{1}}^{1}\right)_{2}$$\scriptstyle\left(J_{x_{1}}^{2}\right)_{2}$$\scriptstyle\left(I_{x_{1}}^{2}\right)_{2}$$\scriptstyle\left(J_{x_{2}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{2}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{1}}^{3}\right)_{2}$$\scriptstyle\left(J_{x_{1}}^{3}\right)_{2}$$\scriptstyle\left(I_{x_{2}}^{2}\right)_{2}$$\scriptstyle\left(J_{x_{2}}^{2}\right)_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$
Figure 1: Here we have shown how to begin constructing $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ with $w=x_{1}x_{1}x_{2}x_{1}^{-1}x_{2}^{-1}$ and $k=2.$ There are $202kl(w)=$ 20 vertices, split in to $k=2$ distinct subsets, each containing $102l(w)=$ 10 vertices each. The vertices on the left (ie with the outer subscript “$1$”) are the $1^{\mathrm{{}^{\prime}st}}$ $w$–loop and the vertices on the right are the $2^{\mathrm{{}^{\prime}nd}}$ $w$–loop. There are $k|w|_{x_{1}}=6$ directed $x_{1}$–edges and $k|w|_{x_{2}}=4$ directed $x_{2}$–edges.
For each $f\in\{x_{1},\dots,x_{r}\}$, we add an undirected $\sigma_{f}$–edge between any two vertices $\left(J_{f}^{j}\right)_{i}$ and $\left(I_{f}^{j^{\prime}}\right)_{i^{\prime}}$ whenever $\sigma_{f}$ dictates that the corresponding indices must be equal in order for $\delta_{\sigma_{f}}\left(J_{f}^{1}\sqcup J_{f}^{2}\sqcup\dots\sqcup J_{f}^{|w|_{f}}\right)$ to be non–zero. This is illustrated in Figure 2.
$\scriptstyle\left(J_{x_{1}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{1}\right)_{1}$$\scriptstyle\left(J_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(J_{x_{2}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{2}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{3}\right)_{1}$$\scriptstyle\left(J_{x_{1}}^{3}\right)_{1}$$\scriptstyle\left(I_{x_{2}}^{2}\right)_{1}$$\scriptstyle\left(J_{x_{2}}^{2}\right)_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle\left(J_{x_{1}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{1}}^{1}\right)_{2}$$\scriptstyle\left(J_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{2}\right)_{2}$$\scriptstyle\left(J_{x_{2}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{2}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{1}}^{3}\right)_{2}$$\scriptstyle\left(J_{x_{1}}^{3}\right)_{2}$$\scriptstyle\left(I_{x_{2}}^{2}\right)_{2}$$\scriptstyle\left(J_{x_{2}}^{2}\right)_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$
Figure 2: Here we continue the construction of the graph from Figure 1 by adding in the $\sigma_{x_{1}}$–edges (in red) and the $\sigma_{x_{2}}$–edges (in purple). In this example, we have $\sigma_{x_{1}}=\big{\{}\{1,3,6\},\ \{2,4,5\}\big{\}}\in\mathrm{Part}\left([3k]\right]$ and $\sigma_{x_{2}}=\big{\{}\{1,3\},\ \{2,4\}\big{\}}\in\mathrm{Part}\left([2k]\right)$.
For each $f\in\{x_{1},\dots,x_{r}\}$, we add undirected $\tau_{f}$–edges similarly, illustrated in Figure 4.
Finally, for each $i\in\{1,\dots,l(w)\},$ we add an undirected $\pi_{i}$–edge between any two vertices whenever $\pi_{i}$ dictates that the corresponding indices must be equal in order for the corresponding inner product term in Definition 4.6 to be non–zero.
In short, for each collection of partitions $\sigma_{f},\tau_{f},\pi_{1},\dots,\pi_{l(w)},$ we draw an undirected edge between any two vertices for which the indices with the same label are necessarily equal for the conditions in Definition 4.6 to hold.
Figure 4 shows a complete example of $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$, with each $\sigma_{f},$ $\tau_{f}$ and $\pi_{i}$ as described in the examples throughout this section.
Definition 4.9.
We will write $\hat{G}\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ for the subgraph of $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ consisting of every vertex and only the undirected edges.
$\scriptstyle\left(J_{x_{1}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{1}\right)_{1}$$\scriptstyle\left(J_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(J_{x_{2}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{2}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{3}\right)_{1}$$\scriptstyle\left(J_{x_{1}}^{3}\right)_{1}$$\scriptstyle\left(I_{x_{2}}^{2}\right)_{1}$$\scriptstyle\left(J_{x_{2}}^{2}\right)_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle\left(J_{x_{1}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{1}}^{1}\right)_{2}$$\scriptstyle\left(J_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{2}\right)_{2}$$\scriptstyle\left(J_{x_{2}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{2}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{1}}^{3}\right)_{2}$$\scriptstyle\left(J_{x_{1}}^{3}\right)_{2}$$\scriptstyle\left(I_{x_{2}}^{2}\right)_{2}$$\scriptstyle\left(J_{x_{2}}^{2}\right)_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$
Figure 3: We continue the construction from Figures 1 and 2. We have added in the $\tau_{x_{1}}$–edges (in dark blue) and the $\tau_{x_{2}}$–edges (in light blue). In this example, $\tau_{x_{1}}=\big{\{}\{1,3,5\}\ \{2,4,6\}\big{\}}$ and $\tau_{x_{2}}=\big{\{}\{1,4\}\ \{2,3\}\big{\}}$.
$\scriptstyle\left(J_{x_{1}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{1}\right)_{1}$$\scriptstyle\left(J_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(J_{x_{2}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{2}}^{1}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{3}\right)_{1}$$\scriptstyle\left(J_{x_{1}}^{3}\right)_{1}$$\scriptstyle\left(I_{x_{2}}^{2}\right)_{1}$$\scriptstyle\left(J_{x_{2}}^{2}\right)_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle\left(J_{x_{1}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{1}}^{1}\right)_{2}$$\scriptstyle\left(J_{x_{1}}^{2}\right)_{1}$$\scriptstyle\left(I_{x_{1}}^{2}\right)_{2}$$\scriptstyle\left(J_{x_{2}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{2}}^{1}\right)_{2}$$\scriptstyle\left(I_{x_{1}}^{3}\right)_{2}$$\scriptstyle\left(J_{x_{1}}^{3}\right)_{2}$$\scriptstyle\left(I_{x_{2}}^{2}\right)_{2}$$\scriptstyle\left(J_{x_{2}}^{2}\right)_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$
Figure 4: This depicts $G\left(\sigma_{f},\tau_{f},\pi_{i}\right),$ where $\sigma_{f}$ and $\tau_{f}$ are as described in Figures 2 and 4 and the $\pi_{i}$ are as follows: $\pi_{1}=\big{\{}\{1\},\ \{2\},\ \{3\}\ \{4\}\big{\}},$ $\pi_{2}=\big{\{}\{1,3\},\ \{2,4\}\big{\}}$, $\pi_{3}=\big{\{}\{1,4\},\ \{2\},\ \{3\}\big{\}}$, $\pi_{4}=\big{\{}\{1\},\ \{3\},\ \{2,4\}\big{\}}$ and $\pi_{5}=\big{\{}\{1,4\},\ \{2,3\}\big{\}}$.
4.5 Obtaining the bound
In the case of $\sigma_{f}=\tau_{f}$ and with all $\pi_{1},\dots,\pi_{l(w)}$ corresponding to permutations (ie with $\sum_{i}\mathrm{del}(\pi_{i})=0$), we can apply [LouderWilton, Theorem 1.2] to obtain the bound of $n^{-k}.$ For the other cases, we need to prove a variant of this theorem that applies in our case. This section is dedicated to this task.
4.5.1 Stackings
Given graph morphisms $\rho_{1}:\Gamma_{1}\to G$ and $\rho_{2}:\Gamma_{2}\to G$, the fibre product $\Gamma_{1}\times_{G}\Gamma_{2}$ is the graph with vertex set
$V\left(\Gamma_{1}\times_{G}\Gamma_{2}\right)=\big{\{}(v_{1},v_{2})\in V\left(\Gamma_{1}\right)\times V\left(\Gamma_{2}\right):\ \rho_{1}\left(v_{1}\right)=\rho_{2}\left(v_{2}\right)\big{\}}$
and edge set
$E\left(\Gamma_{1}\times_{G}\Gamma_{2}\right)=\big{\{}(e_{1},e_{2})\in E\left(\Gamma_{1}\right)\times E\left(\Gamma_{2}\right):\ \rho_{1}\left(e_{1}\right)=\rho_{2}\left(e_{2}\right)\big{\}},$
where $t\left(e_{1},e_{2}\right)=\left(t\left(e_{1}\right),t\left(e_{2}\right)\right)$ and $h\left(e_{1},e_{2}\right)=\left(h\left(e_{1}\right),h\left(e_{2}\right)\right).$
Louder and Wilton [LouderWilton] first developed the notion of a stacking of a graph immersion and we adapt their definitions slightly to suit our case.
Definition 4.10.
Let $\Gamma$ be a finite graph and $\mathbb{S}$ a $1$–complex with an immersion $\Lambda:\mathbb{S}\to\Gamma.$ A stacking is an embedding $\hat{\Lambda}:\mathbb{S}\to\Gamma\times\mathbb{R}$ such that $\pi\hat{\Lambda}=\Lambda,$ where $\pi:\Gamma\times\mathbb{R}\to\Gamma$ is the trivial $\mathbb{R}$–bundle.
Let $\eta:\Gamma\times\mathbb{R}\to\mathbb{R}$ be the projection to $\mathbb{R}.$
Definition 4.11.
Given a stacking $\hat{\Lambda}:\mathbb{S}\to\Gamma\times\mathbb{R}$ of an immersion $\Lambda:\mathbb{S}\to\Gamma,$ define
$\mathcal{A}_{\hat{\Lambda}}\overset{\mathrm{def}}{=}\big{\{}x\in\mathbb{S}:\ \forall y\neq x,\ \mathrm{if}\ \Lambda(x)=\Lambda(y)\ \mathrm{then}\ \eta\left(\hat{\Lambda}(x)\right)>\eta\left(\hat{\Lambda}(y)\right)\big{\}}$
and
$\mathcal{B}_{\hat{\Lambda}}\overset{\mathrm{def}}{=}\big{\{}x\in\mathbb{S}:\ \forall y\neq x,\ \mathrm{if}\ \Lambda(x)=\Lambda(y)\ \mathrm{then}\ \eta\left(\hat{\Lambda}(x)\right)<\eta\left(\hat{\Lambda}(y)\right)\big{\}}.$
4.5.2 Bounding expected character
By Remark 1.6, we only need to prove the bound in Theorem 1.5 in the case where $w$ is not the identity or a proper power. From now, assume that $w$ is not a proper power.
For each $\sigma_{f},\tau_{f},\pi_{i},$ construct the graph $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ from Section 4.4, consisting of $kl(w)$ disjoint, directed $f$ edges organised into $k$ $w$–loops and then marking on the blocks of the partitions $\sigma_{f},\tau_{f},\pi_{i}$ using undirected edges.
We construct a new graph $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ by:
  • gluing together any vertices that are connected by a $\sigma_{f},\tau_{f}$ or $\pi_{i}$–edge,
  • if $\sigma_{f}$ and $\tau_{f}$ connect both the initial and terminal vertices of some collection of $f$–edges, we merge these in to a single $f$–edge, see figure 5.
Lemma 4.12.
Given $\sigma_{f},\tau_{f}\in\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right)$ and $\pi_{i}\in\mathrm{Part}_{\leq S_{k}}\left([2k]\right),$
$\sum_{f}|\sigma_{f}\wedge\tau_{f}|=\Big{|}E\left(\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)\right)\Big{|}.$
Proof.
Every block of $\sigma_{f}\wedge\tau_{f}$ of size $p$ corresponds to some collection of $f$–edges in $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ of size $p$, whose initial and terminal vertices have been glued together and whose edges have been merged in the construction of $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right).$ Every $f$–edge in $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ then corresponds to a block of $\sigma_{f}\wedge\tau_{f}.$
Lemma 4.13.
Given $\sigma_{f},\tau_{f}\in\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right)$ and $\pi_{i}\in\mathrm{Part}_{\leq S_{k}}\left([2k]\right),$
$\mathcal{N}(\sigma_{f},\tau_{f},\pi_{i})\ll n^{\left|V\left(\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)\right)\right|}.$
Proof.
Each vertex of $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ corresponds to a connected component of $\hat{G}\left(\sigma_{f},\tau_{f},\pi_{i}\right)$. Any collection of indices must be equal if their corresponding vertices are in the same connected component of $\hat{G}\left(\sigma_{f},\tau_{f},\pi_{i}\right)$. Clearly, if there were no more restrictions, we would have
$\mathcal{N}(\sigma_{f},\tau_{f},\pi_{i})=n^{\left|V\left(\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)\right)\right|}.$
The lemma follows since we have the additional restriction that within each multi–index, each index must be distinct. ∎
So (24) becomes
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]\ll_{k,l(w)}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}n^{-\sum_{i}\mathrm{del}(\pi_{i})}n^{\chi\left(\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)\right)}.$ (25)
$\scriptstyle\left[I_{x_{2}}^{2}\right]_{1}=\scriptstyle\left[I_{x_{2}}^{1}\right]_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$$\scriptstyle x_{2}$$\scriptstyle x_{1}$
Figure 5: Above we show how to construct $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ from the graph $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ constructed in Figure 4. We have labeled one vertex to show which vertices of $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ have been glued together in the construction.
Remark 4.14.
It is worth noting here why the most obvious approach to obtaining the required bound does not work. We can ‘complete’ each $\pi_{i}$ to a permutation $\hat{\pi}_{i}$ by gluing together vertices of $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ according to the additional identifications dictated by the edges that must be added to $\pi_{i}$ to construct $\hat{\pi}_{i}.$ Then, one may observe that
$n^{-\sum_{i}\mathrm{del}(\pi_{i})}n^{\chi\left(\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)\right)}\ll n^{\chi\left(\Gamma\left(\sigma_{f},\tau_{f},\hat{\pi}_{i}\right)\right)}$
and seek to apply [LouderWilton, Theorem 1.2] directly. Unfortunately, in gluing the vertices of $\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ together, one may lose the distinctness property that no two vertices of $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ labeled, say, $\left(J_{f}^{i}\right)_{j_{1}}$ and $\left(J_{f}^{i}\right)_{j_{2}}$ with $j_{1}\neq j_{2}$, can be glued together, which would then weaken the bound achieved by applying [LouderWilton, Theorem 1.2]. Therefore, we proceed with our extension of their result.
Define $\mathcal{B}_{r}\overset{\mathrm{def}}{=}\bigvee_{i=1}^{r}S_{i}^{1},$ the bouquet of $r$ oriented circles, labeled $x_{1},\dots,x_{r}$ and with wedge point labeled $o$. Define $W$ to be the graph consisting of a single cycle of oriented $f$–labeled edges that, when traversed in the correct direction, reads out the word $w.$ Then $W$ comes equipped with an obvious primitive (since $w$ is not a proper power) immersion
$\Lambda:W\to\mathcal{B}_{r}.$
By [LouderWilton, Lemma 3.4], this immersion has a stacking
$\hat{\Lambda}:W\to\mathcal{B}_{r}\times\mathbb{R}.$
For each $\sigma_{f},\tau_{f},\pi_{i},$ write $\Gamma=\Gamma\left(\sigma_{f},\tau_{f},\pi_{i}\right).$ We have a map
$\rho:\Gamma\to\mathcal{B}_{r}$
and we can form the fibre product graph
$\Gamma\times_{\mathcal{B}_{r}}W.$
We denote by $\mathbb{S}$ the components of the fibre product graph defined by following the paths in $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ which alternate between $f$–edges and $\pi_{i}$–edges. This defines some partition of $w^{k}.$
The connected components of $\mathbb{S}$ are either circles that read out $w^{k_{i}}$ or pieces, which consist of two closed endpoints and a chain of vertices of valence two in between them. Each piece reads out some subword of $w^{k}.$ The concatenation of the words spelled out by all the connected components of $\mathbb{S}$ is exactly $w^{k}$ and the number of pieces of $\mathbb{S}$ is exactly $\sum_{i}\mathrm{del}\left(\pi_{i}\right).$
We have an immersion
$\Lambda^{\prime}:\mathbb{S}\to\Gamma$
and an immersion
$\delta:\mathbb{S}\to W.$
By the same argument as [LouderWilton, Lemma 2.5], we have a stacking,
$\hat{\Lambda}^{\prime}:\mathbb{S}\to\Gamma\times\mathbb{R}.$
Lemma 4.15.
Every vertex in $\Gamma$ is covered at least twice by $\Lambda^{\prime}:\mathbb{S}\to\Gamma.$
Proof.
Every vertex $v$ of $\Gamma$ is constructed by gluing together at least two vertices in $G\left(\sigma_{f},\tau_{f},\pi_{i}\right),$ and $v$ is covered by $\Lambda^{\prime}$ exactly once for each vertex in $G\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ that has been glued in the construction of $v$. ∎
The connected components of $\mathbb{S}$ are then one of the following:
  • a circle
  • a closed arc: a connected and simply connected union of vertices and interiors of edges, with both ends closed (ie, a piece).
If we also define the following:
  • an open arc: a connected and simply connected union of vertices and interiors of edges, with both ends open
  • a half–open arc: a connected and simply connected union of vertices and interiors of edges, with one end open and one end closed,
then the following lemma is immediate. [^4] We are especially indebted to Noam Ta Shma for pointing out an error in our original argument for proving the extension of Wise’s $w$–cycles conjecture and for his feedback on the correction presented from this point onward.
Lemma 4.16.
Every connected component of $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ or $\mathcal{B}_{\hat{\Lambda}^{\prime}}$ is either a circle, an open arc, a half–open arc or a closed arc.
Lemma 4.17.
Neither $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ or $\mathcal{B}_{\hat{\Lambda}^{\prime}}$ contain a circle, a closed arc reading out a word of length $\geq l(w)$ or an open/half–open arc reading out a word of length $>l(w).$
Proof.
Let $C\subseteq\mathbb{S}$ be a circle, a closed arc reading out a word of length $\geq l(w)$, or an open/half–open arc reading out a word of length $>l(w),$ which is contained inside $\mathcal{A}_{\hat{\Lambda}^{\prime}}$. Then the map
$\delta\big{|}_{C}:C\to W$
is surjective.
$W$ must contain a vertex $x$ belonging to $\mathcal{B}_{\hat{\Lambda}}.$ Then, the preimage $\left(\delta\big{|}_{C}\right)^{-1}(x)$ contains a vertex $x^{\prime}$ belonging to $\mathcal{B}_{\hat{\Lambda}^{\prime}}$, so that $x^{\prime}\in\mathcal{A}_{\hat{\Lambda}^{\prime}}\cap\mathcal{B}_{\hat{\Lambda}^{\prime}}.$ This implies that $\Lambda^{\prime}(x^{\prime})\in V(\Gamma)$ is covered exactly once by $\Lambda^{\prime},$ contradicting Lemma 4.15. To prove the lemma for $\mathcal{B}_{\hat{\Lambda}^{\prime}},$ we can swap $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ and $\mathcal{B}_{\hat{\Lambda}^{\prime}}$ in the proof. ∎
So the connected components of $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ are one of the following:
  • open arcs that read out words of length $\leq l(w)$,
  • half–open arcs that read out words of length $\leq l(w)$ or
  • closed arcs that read out words of length $<l(w).$
The main points here are that every connected component has two ends (that are either open or closed) and that none of the connected components can be too long. Moreover, we observe that the only closed ends that can be in $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ or $\mathcal{B}_{\hat{\Lambda}^{\prime}}$ are those that are the closed ends of pieces in $\mathbb{S}.$
Write
$\mathcal{A}\overset{\mathrm{def}}{=}\#\big{\{}\mathrm{ends\ of\ connected\ components\ of}\ \mathcal{A}_{\hat{\Lambda}^{\prime}}\big{\}}$
and
=$\mathcal{B}\overset{\mathrm{def}}{=}\#\big{\{}\mathrm{ends\ of\ connected\ components\ of}\ \mathcal{B}_{\hat{\Lambda}^{\prime}}\big{\}}.$
Lemma 4.18.
We have $\mathcal{A}\geq 2k$ and $\mathcal{B}\geq 2k.$
Proof.
Pick any vertex $w\in W$ such that $w\in\mathcal{A}_{\hat{\Lambda}}.$ Then $w$ has at least $k$ preimages $\delta^{-1}(w)$ in $\mathbb{S},$ all of which must be contained in $\mathcal{A}_{\hat{\Lambda}^{\prime}}.$ If two such preimages belong to the same connected component $C$ of $\mathcal{A}_{\hat{\Lambda}^{\prime}},$ then $C$ must either be a circle, a closed arc reading out a word of length $\geq l(w)$ or an open/half–open arc reading out a word of length $>l(w),$ contradicting Lemma 4.17. Thus, there must be at least $k$ connected components in $\mathcal{A}_{\hat{\Lambda}^{\prime}},$ so there must be at least $2k$ ends.
Repeating the argument with $w\in\mathcal{B}_{\hat{\Lambda}}$ proves the lemma for $\mathcal{B}.$
Write
$\mathcal{A}_{\mathrm{closed}}\overset{\mathrm{def}}{=}\#\big{\{}\mathrm{closed\ ends\ of\ connected\ components\ of}\ \mathcal{A}_{\hat{\Lambda}^{\prime}}\big{\}}$
and
$\mathcal{B}_{\mathrm{closed}}\overset{\mathrm{def}}{=}\big{\{}\mathrm{closed\ ends\ of\ connected\ components\ of}\ \mathcal{B}_{\hat{\Lambda}^{\prime}}\big{\}}.$
Lemma 4.19.
We have
$\mathcal{A}=-2\chi+2\mathcal{A}_{\mathrm{closed}}$
and
$\mathcal{\mathcal{B}}=-2\chi+2\mathcal{B}_{\mathrm{closed}}.$
Proof.
We prove the lemma for $\mathcal{A}_{\hat{\Lambda}^{\prime}}.$ Each vertex $v\in V(\Gamma)$ is covered by $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ once. If $v$ is covered by a connected component of $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ passing through one of $v$’s preimages in $\mathbb{S},$ then there must be
$\deg(v)-2$
open ends of $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ that end at the other preimages of $v$ in $\mathbb{S},$ to cover the other edges of $\Gamma$ that are incident at $v.$
Otherwise, $v$ is covered by $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ by a closed end of a connected component of $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ that ends at one of $v$’s preimages. In this case, we count exactly $1$ closed end of $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ that ends at one of $v$’s preimages and exactly
$\deg(v)-1$
open ends of $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ that end at the other preimages of $v$ in $\mathbb{S},$ to cover the other edges of $\Gamma$ that are incident at $v$.
In total, we see that
$\mathcal{A}=\sum_{v}(\deg(v)-2)+\sum_{u}\deg(u),$
where the first sum is over $v\in V(\Gamma)$ that are covered by $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ by a connected component passing through one of their preimages and the second sum is over $u\in V(\Gamma)$ that are covered by $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ by a closed end at one of their preimages. This can be rewritten as
$\sum_{v\in V(\Gamma)}(\deg(v)-2)+2\sum_{u}=-2\chi(\Gamma)+2\mathcal{A}_{\mathrm{closed}}.$
Combining Lemma 4.18 with Lemma 4.19 implies that
$k\leq-\chi(\Gamma)+\mathcal{A}_{\mathrm{closed}}$
and also that
$k\leq-\chi(\Gamma)+\mathcal{B}_{\mathrm{closed}}.$
This implies that
$2k\leq-2\chi(\Gamma)+\mathcal{A}_{\mathrm{closed}}+\mathcal{B}_{\mathrm{closed}}\leq-2\chi(\Gamma)+2\sum_{i}\mathrm{del\left(\pi_{i}\right),}$ (26)
since the total number of closed ends in $\mathbb{S}$ is $2\sum_{i}\mathrm{del}\left(\pi_{i}\right)$ and each closed end in $\mathbb{S}$ can be, at most, a closed end in $\mathcal{A}_{\hat{\Lambda}^{\prime}}$ or a closed end in $\mathcal{B}_{\hat{\Lambda}^{\prime}},$ but not both, since that would imply that its image in $\Gamma$ is covered only once by the immersion $\Lambda^{\prime}:\mathbb{S}\to\Gamma$, contradicting Lemma 4.15.
Combining (26) with (25) proves Theorem 1.5.
4.6 Proof of Theorem 4.1
Let $w\in F_{r}=\left\langle x_{1},\dots,x_{r}\right\rangle$ for $r$ fixed. Let $m\leq r$ be the minimum number of generators $x_{i_{1}},\dots,x_{i_{m}}$ such that $w$ can be written using the alphabet $\{x_{i_{1}},x_{i_{1}}^{-1},\dots,x_{i_{m}},x_{i_{m}}^{-1}\}$. So, up to relabeling the generators, $w\in F_{m}=\left\langle x_{1},\dots,x_{m}\right\rangle,$ with $m\leq r.$
Beginning with (24), we have
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$
$=$ $d_{\lambda}\sum_{\sigma_{f},\tau_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{i=1}^{l(w)}c(n,k,\lambda,\pi_{i})\right)\left(\prod_{f\in\{x_{1},\dots,x_{m}\}}\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)\right)$
$\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right).$
For each $f\in\{x_{1},\dots,x_{m}\},$ for each $\sigma_{f}\in\overset{\star}{\mathrm{Part}}\left(\left[|w|_{f}k\right]\right),$
$\mathrm{Wg}_{n,\left(|w|_{f}k\right)}\left(\sigma_{f},\tau_{f}\right)=\sum_{\rho_{f}\leq\sigma_{f}\wedge\tau_{f}}\mu(\rho_{f},\sigma_{f})\mu(\rho_{f},\tau_{f})\frac{1}{(n)_{|\rho_{f}|}}.$
This is equal to
$\frac{C_{1}}{(n)_{|w|_{f}k}}+\frac{C_{2}}{(n)_{|w_{f}|k-1}}+\dots+\frac{C_{|w|_{f}k-|\sigma_{f}\wedge\tau_{f}|+1}}{(n)_{|\sigma_{f}\wedge\tau_{f}|}}$
$=$ $\frac{g_{\sigma_{f},\tau_{f}}(n)}{(n)_{|w|_{f}k}},$
where $g_{\sigma_{f},\tau_{f}}(n)$ is a polynomial in $n$ of maximum degree $|w|_{f}k-|\sigma_{f}\wedge\tau_{f}|$.
For each $\pi\in\mathrm{Part}_{\leq S_{k}}\left([2k]\right)$, we have
$c(n,k,\lambda,\pi)=\frac{(-1)^{k}d_{\lambda}^{2}(n)_{2k}}{\left(k!\right)^{2}(n)_{\lambda}}\left[\sum_{\sigma,\sigma^{\prime}\in S_{k}}\sum_{\pi^{\prime}\leq\iota\left(\sigma^{-1}\sigma^{\prime}\right)\wedge\pi}\frac{\mu\left(\pi^{\prime},\pi\right)(-1)^{|\pi^{\prime}|}\chi^{\lambda}\left(\sigma\right)\chi^{\lambda}\left(\sigma^{\prime}\right)}{(n)_{|\pi^{\prime}|}}\right],$
where $(n)_{\lambda}=\prod_{j=1}^{k}\left(n-k+1+\check{\lambda}_{j}-j\right).$ The term inside the square bracket is equal to
$\frac{\underset{\sigma,\sigma^{\prime}\in S_{k}}{\sum}\chi^{\lambda}(\sigma)\chi^{\lambda}\left(\sigma^{\prime}\right)+\dots+\kappa_{\pi}\left(n-|\pi|\right)\dots\left(n-2k+1\right)}{(n)_{2k}},$
with
$\kappa_{\pi}=\left(\underset{\sigma,\sigma^{\prime}\ \mathrm{s.t.}\ \pi\leq\iota\left(\sigma^{-1}\sigma^{\prime}\right)}{\sum}(-1)^{|\pi|}\chi^{\lambda}\left(\sigma\right)\chi^{\lambda}\left(\sigma^{\prime}\right)\right).$
So then $c(n,k,\lambda,\pi)$ is equal to
$\frac{h_{k,\lambda,\pi}(n)}{(n)_{\lambda}},$
where
$h_{k,\lambda,\pi}=\frac{(-1)^{k}d_{\lambda}^{2}}{\left(k!\right)^{2}}\Bigg{[}\underset{\sigma,\sigma^{\prime}\in S_{k}}{\sum}\chi^{\lambda}(\sigma)\chi^{\lambda}\left(\sigma^{\prime}\right)+\kappa_{\pi}\left(n-|\pi|\right)\dots\left(n-2k+1\right)\Bigg{]},$
a polynomial in $n$ of maximum degree $2k-|\pi|=k-\mathrm{del}(\pi).$
It follows that
$\prod_{i=1}^{l(w)}c(n,k,\lambda,\pi_{i})=\frac{\prod_{i=1}^{l(w)}h_{k,\lambda,\pi_{i}}(n)}{\left(n\right)_{\lambda}^{l(w)}}.$
Hence,
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=\frac{p_{k,\lambda,w}\left(n\right)}{\left(n\right)_{\lambda}^{l(w)}\prod_{f}(n)_{|w|_{f}k}},$ (27)
where
$p_{k,\lambda,w}\left(n\right)=d_{\lambda}\sum_{\sigma_{f}}^{\star}\sum_{\pi_{1},\dots,\pi_{l(w)}}^{\leq S_{k}}\left(\prod_{f}g_{\sigma_{f},\tau_{f}}(n)\right)\left(\prod_{i=1}^{l(w)}h_{k,\lambda,\pi_{i}}(n)\right)\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right).$
For each collection of $\sigma_{f},\tau_{f},\pi_{i}$:
  • $\left(\prod_{f}g_{f}(n)\right)$ has maximum degree
    $\sum_{f}\left(|w|_{f}k-|\sigma_{f}\wedge\tau_{f}|\right)=kl(w)-\sum_{f}|\sigma_{f}\wedge\tau_{f}|,$
  • $\left(\prod_{i=1}^{l(w)}h_{k,\lambda,\pi_{i}}(n)\right)$ has maximum degree
    $\sum_{i=1}^{l(w)}k-\mathrm{del}(\pi_{i})=kl(w)-\sum_{i}\mathrm{del(\pi_{i})},$
  • $\mathcal{N}\left(\sigma_{f},\tau_{f},\pi_{i}\right)$ has maximum degree
    $-k+\sum_{i}\mathrm{del}(\pi_{i})+\sum_{f}|\sigma_{f}\wedge\tau_{f}|,$
so that the degree of $p_{k,\lambda,w}(n)$ is less than or equal to
$2kl(w)-k.$
We can rewrite the denominator of (27) using reciprocal polynomials. We have
$\prod_{f}(n)_{|w|_{f}k}$
$=$ $\prod_{d=1}^{m}\prod_{c=1}^{|w|_{x_{d}}k-1}(n-c)$
$=$ $n^{kl(w)}\prod_{d=1}^{m}\prod_{c=1}^{|w|_{x_{d}}k-1}\left(1-c.\frac{1}{n}\right).$
Similarly,
$\left(n\right)_{\lambda}^{l(w)}=n^{kl(w)}\left[\prod_{j=1}^{k}\left(1+\left(1+\check{\lambda}_{j}-j-k\right).\frac{1}{n}\right)\right]^{l(w)}.$
It follows that
$\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]=\frac{\frac{1}{n^{2kl(w)}}p_{k,\lambda,w}\left(n\right)}{\tilde{g}\left(\frac{1}{n}\right)}=\frac{\hat{P}_{w,k,\lambda}\left(\frac{1}{n}\right)}{\tilde{g}\left(\frac{1}{n}\right)},$
where
$\tilde{g}\left(x\right)=\prod_{d=1}^{m}\prod_{c=0}^{|w|_{x_{d}}k-1}\left(1-cx\right)\left[\prod_{j=1}^{k}\left(1+\left(1+\check{\lambda}_{j}-j-k\right)x\right)\right]^{l(w)}.$
The numerator $\hat{P}_{w,k,\lambda}\left(\frac{1}{n}\right)$ is clearly a polynomial in $\frac{1}{n}$ of maximum degree $2kl(w)$.
Now assume that $l(w)\leq q$, which also implies that $m\leq q.$ Then $\tilde{g}(x)$ always divides
$\hat{g}(x)=\prod_{c=1}^{q}\left(1-cx\right)^{q}\left[\prod_{j=1}^{k}\left(1+\left(1+\check{\lambda}_{j}-j-k\right)x\right)\right]^{q}.$
The sequence $\check{\lambda}_{1}-1,\dots,\check{\lambda}_{k}-k$ is strictly decreasing. Moreover, for any $\lambda\vdash k$ and any $j\in[k],$
$1+\check{\lambda}_{j}-j-k\in[1-2k,0],$
so that
$\hat{g}(x)\Biggr{|}\prod_{c=1}^{q}\left(1-cx\right)^{q}\left[\prod_{j=1}^{2k}\left(1+\left(1-j\right)x\right)\right]^{q}=g_{q,k}(x).$
It follows that $\mathbb{E}_{w}\left[\chi^{\lambda^{+}(n)}\right]$ can be written
$\frac{P_{w,k,\lambda}\left(\frac{1}{n}\right)}{g_{q,k}\left(\frac{1}{n}\right)},$
where
[…]
This is a polynomial in $\frac{1}{n}$ of maximum degree
$2kl(w)+k(2q-l(w))+\sum_{d=1}^{m}\left(q-|w|_{x_{d}}k+1\right)$
$\leq$ $2kq+mq+m$
$\leq$ $3kq+q^{2}.$
[…]
Ewan Cassidy, Department of Mathematical Sciences, Durham University, Lower Mountjoy, DH1 3LE, Durham, United Kingdom
ewan.g.cassidy@durham.ac.uk